### v1

# Using Large Language Models for Expert Prior Elicitation in Predictive Modelling

Abstract
Large language models (LLMs), trained on diverse data effectively acquire a breadth of information across various domains. However, their computational complexity, cost, and lack of transparency hinder their direct application for specialised tasks. In fields such as clinical research, acquiring expert annotations or prior knowledge about predictive models is often costly and time-consuming. This study proposes using LLMs to elicit expert prior distributions for predictive models. This approach also provides an alternative to in-context learning, where language models are tasked with making predictions directly. We compare LLM-elicited and uninformative priors, evaluate whether LLMs truthfully generate parameter distributions, and propose a model selection strategy for in-context learning and prior elicitation. Our findings show that LLM-elicited prior parameter distributions significantly reduce predictive error compared to uninformative priors in low-data settings. Applied to clinical problems, this translates to fewer required biological samples, lowering cost and resources. Prior elicitation also consistently outperforms and proves more reliable than in-context learning at a lower cost, making it a preferred alternative in our setting. We demonstrate the utility of this method across various use cases, including clinical applications. For infection prediction, using LLM-elicited priors reduced the number of required labels to achieve the same accuracy as an uninformative prior by 55%, at 200 days earlier in the study.
1 Introduction
In fields such as clinical research, labelling data can be costly and time-consuming [1]. For example, in healthcare monitoring, patients’ data is recorded through wearables or remote monitoring devices and observations are continuously collected. However, the labelling of such data for clinical outcomes, such as diagnosing illnesses or collecting and studying samples, can be challenging. Further, estimating uncertainty on clinical predictions is paramount [2], especially when available training data is sparse.
Expert knowledge about a problem domain can provide prior distributions on model parameters. This can improve a model’s performance when trained with few data points and enable estimates of uncertainties on parameters and predictions. However, expert knowledge about model parameter distributions is often hard to obtain due to the challenges of performing prior elicitation from experts [3, 4].
Due to their diverse training data, language models offer an opportunity to begin resolving these issues. By considering a language model as an “expert” in a given field, we can extract insights about a predictive task that might lead to improved predictive models. This, and the abilities of language models to reason about problem settings outside of their intended use cases [5, 6, 7, 8, 9, 10], have inspired their use in performing prior elicitation [11, 12, 13]. Specifically, Selby et al. [12] show, by considering prior predictive distributions, that GPT-4 [10] can provide parameter distributions that allow for zero-shot performance (making predictions without training examples) of Bayesian models that compare to models trained on tens of data points. Further, in Requeima et al. [13], the authors prompt language models with text and collected data points to elicit posterior predictive distributions for regression tasks, which they find outperforms Gaussian processes with radial basis function kernels on some tasks. This is done by repeatedly sampling a language model at a new feature value, given the previous input and output observations and a given text description that contains prior information from a human expert. Given the rate of development of language models [14, 15, 16, 17], it is conceivable that their knowledge improves and broadens with time, motivating the construction of a probabilistic framework to reason about elicited prior task knowledge from language models, as well as understand their use for predictive tasks.
Figure 1: Prior elicitation using a language model.
In-context learning allows for the direct use of language models for predictive tasks. Here, language models are provided task descriptions and demonstration examples from a training data set (in text form) and are prompted to make predictions on unseen data [18, 19]. In this sense, in-context learning is similar to eliciting a prior distribution on a predictive model’s parameters and calculating its posterior predictive distribution, where the main differences arise from: (1) Language models are notoriously weak at performing arithmetic on numerical values [20]; (2) Language models lack transparency and may not be learning from demonstrations in a predictable or Bayesian way [21]; (3) Language models may not be able to faithfully provide their internal knowledge of a task when specifying a prior parameterisation for a separate predictive model or use that same prior knowledge in-context. To study the trade-offs between using in-context learning and prior elicitation we aim to use Bayes factor for model selection, and propose an approach to estimate a language model’s internal predictive prior and posterior models.
We are motivated by the following questions, and make the following contributions:
  1. 1.
    Question: Can language models provide good prior distributions over the parameters of a predictive model, which enables the training of more robust models with less data?
    Method: We propose a method to elicit general prior parameter distributions for linear models from language models for predictive tasks (Section 2.2).
    Finding: We find that using language models for expert prior elicitation provides better posterior accuracy and mean square error than an uninformative prior on a variety of numerical predictive tasks, enabling the construction of predictive models with fewer labelled examples (Section 3.1).
  2. 2.
    Question: How do the LLM’s in-context prior distributions compare to the ones we elicit?
    Method: We present an approach for extracting the in-context prior distribution (Section 2.3.2).
    Finding: We find these distributions differ significantly from those that we elicit (Section 3.4).
  3. 3.
    Question: How can we check if the language model is performing Bayesian inference in-context?
    Method: By additionally providing the language model with training demonstrations, we additionally extract the LLM’s in-context predictive parameter posterior distribution (Section 2.3.2). We can then perform Monte Carlo sampling on the extracted prior and measure its difference to the extracted posterior.
    Finding: We find that the language is inconsistent in its approximation of Bayesian inference, and that its reliability depends on the task (Section 3.5).
  4. 4.
    Question: When is the performance better with in-context learning?
    Method: We use the Bayes factor as a model selection method to decide between prior elicitation and in-context learning given a task (Section 2.4).
    Finding: Before considering cost and reliability, the Bayes factor depends on the task. However, after considering the in-context model’s inconsistent approximation of Bayesian inference and its cost, we conclude that prior elicitation is the preferred choice for all tasks tested (Section 3.6).
For implementing our methods and reproducing our results, we provide details about the code [^1] https://github.com/alexcapstick/llm-elicited-priors in Appendix A.1 and complete experimental information in Appendix A.5.
2 Methods
2.1 Defining the task descriptions
Our goal in this section is to define the space of prompts that we use to elicit priors and gain predictions from a language model. We start by defining the space of all tasks as $\mathcal{T}$. An individual task, such as a specific classification or regression problem, will be referred to as $T\in\mathcal{T}$. Since a description of a task can be rephrased many times without changing its meaning, we need to consider a space of possible task descriptions, referred to as $\mathcal{I}$, for each task. We say that the probability of a random text sequence $I\in\mathcal{I}$ being a relevant task description for task $T\in\mathcal{T}$ is: $p(I|T)$.
We assume that all relevant task descriptions have equal probability, i.e. if $I_{i}$ and $I_{j}$ fully describe $T$, then $p(I_{i}|T)=p(I_{j}|T)$. We also assume that irrelevant task descriptions have $p(I|T)=0$. This means that for a task $T$, all of the mass of $p(\mathcal{I}|T)$ is uniformly spread across the relevant task descriptions only.
In reality, the number of relevant task descriptions would be too large to cover since each word in the task description could be replaced with a synonym without losing the meaning of the full description. In Appendix A.11, we show that there are diminishing returns as you increase the number of ways of describing a task. For our experiments, we use $100$ relevant task descriptions to cover this space.
In practice, to produce $100$ task descriptions, we employ a language model to rephrase a single human written system role and user role $10$ times each and take their product. This means that using $100$ task descriptions is not particularly time-consuming and could be increased significantly by providing many base descriptions or prompting the language model to produce many more variants.
2.2 Using language models to elicit expert priors
2.2.1 Background
Bayesian inference allows us to incorporate our knowledge about a predictive task before observing data. This is done through the use of the prior knowledge $p(\theta)$ (over a predictive model parameterised by $\theta$), which can be our own intuition about which features are more important for a prediction and in which way they correlate. Additionally, using a prior allows us to report uncertainty over new predictions $(\tilde{x},\tilde{y})$, through the posterior predictive distribution $p(y|\tilde{x},D)$. If the prior is well-specified, using a prior can also reduce the number of data points required to construct a comparably accurate predictive model.
A key problem in machine learning is defining this prior distribution $p(\theta)$ given expert knowledge. Previous work studied how this distribution can be elicited from experts [3], which can be challenging and costly. In parallel, research has shown that large language models (LLMs) encode knowledge about a surprising variety of tasks [5, 6, 7, 8, 9, 10], which could be used to generate information about a predictive task [11, 12, 13], possibly supplemented with scientific discussions [22].
Choi et al. [11] explores the use of an LLM to perform feature selection, causal discovery, and to modify the reward function in a reinforcement learning task. In the first case, they observe that when listing features from two datasets, GPT-3 can correctly identify the features that correspond to a single specified predictive task. Additionally, when applied to causal discovery, the language model is often able to predict the directionality of causal relationships in the Tüebingen Cause-Effect Pairs dataset [23]. This suggests the language model’s utility in providing a prior distribution over the parameters of a predictive model, inspiring us to explore the possibility of using LLMs for eliciting prior parameter distributions and allowing a direct comparison to in-context learning.
Further, Selby et al. [12] elicit parameters of a beta distribution to define a prior for predictive tasks where results from human experts are available. The priors from four LLMs were visually compared to those elicited from human experts using the SHELF [3] method, which demonstrated significant differences between experts and LLMs. In our work, we extend this by using the LLM to elicit multiple Gaussian distributions as priors. We combine these to create a mixture model for a more general distribution over our parameters and test their utility for linear predictive models.
Additionally, Zhu and Griffiths [24] describe using iterated learning to elicit prior distributions over language models for causal learning, proportion estimation, and predicting everyday quantities. The authors employ a Markov chain Monte Carlo method to access an LLM’s prior knowledge by performing successive inference using previous model outputs that tend towards a limiting distribution. By doing this, the authors assume that in-context learning is a form of Bayesian inference. However, in their work Zhu and Griffiths [24] do not consider how prior elicitation from LLMs can be used for predictive models or test whether an LLM is performing Bayesian inference, both of which we study in this work.
2.2.2 Our contribution: LLM-elicited priors
Our first contribution, therefore, focuses on proposing a method for eliciting an LLM’s prior knowledge of a task to construct general distributions on the parameters of a predictive model that improve performance with less data compared to a predictive model with an uninformative prior and produce uncertainty about its predictions. An overview of this process is given in Figure 1.
Formally, given a language model $M$ and task $T$, we will obtain a single Gaussian prior for each feature, for each task description $I$. This is done by sampling $(\mu_{k},\sigma_{k})$ from $p_{M,T}(\mu,\sigma|I_{k})$, where we obtain a single $\mu_{k}$ and $\sigma_{k}$ by prompting the language model to provide a guess of the mean and standard deviation of a Gaussian prior for each feature on a linear model.
By taking a mixture of these individual Gaussian distributions, we construct a prior $p(\theta|M,T)=p_{M,T}(\theta)$ over our model parameters $\theta$:
$p_{M,T}(\theta)=\sum_{k=1}^{K}\pi_{k}\mathcal{N}(\theta|\mu_{k},{\sigma_{k}}^{2})\\ ~{}~{}\text{where}~{}(\mu_{k},\sigma_{k})\sim p_{M,T}(\mu,\sigma|I_{k})~{}\text{and}~{}\pi_{k}\sim\text{Dir}(1)$ (1)
In our experiments, our mixture is over $100$ components, using several different task descriptions. Mixtures of Gaussians are able to approximate any probability density with enough components, allowing us to approximate the language model’s belief about the task, even if non-Gaussian. We would also expect the language model to have learnt how to parameterise a Gaussian distribution from its training, making them good candidates. However, mixtures of other distributions can be used instead and may be more useful for particular tasks.
The posterior predictive distribution on unseen data $(\tilde{x},\tilde{y})$, given an observed dataset $\mathcal{D}$, becomes:
$p_{M,T}(y|\tilde{x},\mathcal{D})=\int_{\Theta}\sum_{k=1}^{K}p_{M,T}(y|\tilde{x},\theta)p_{M,T}(\theta|\mathcal{D},I_{k})p_{M,T}(I_{k})d\theta$
Here, $p_{M,T}(I)$ refers to the probability distribution discussed in Section 2.1. The distribution $p_{M,T}(\theta|\mathcal{D},I)$ describes the posterior distribution over the parameterisation of a predictive model, which is calculated using the available data $\mathcal{D}$ and the prior parameterisation we extract from the language model.
The practical sampling of these distributions is discussed in Section 3.1 and Appendix A.5.2.
Since we are using a language model’s prior knowledge and a given training dataset to make predictions, this method is an alternative to in-context learning. The difference is that prior elicitation uses a marginal distribution and Bayesian inference to address some of the shortfalls of in-context learning.
2.3 In-context learning
2.3.1 Background
In-context learning aims to use the knowledge that language models acquire during training to make predictions on unseen data. Specifically, this method requires providing a language model with a description of the predictive task and some demonstration examples that allow it to learn “in-context” before the user queries the predictions on some unseen data.
Probabilistically, given a language model $M$ and task $T$, and many task descriptions $I_{k}\in\mathcal{I}$, a prediction of $\tilde{y}$ on features $\tilde{x}$ using in-context learning with observed data $\mathcal{D}$ is:
$p_{M,T}(y|\tilde{x},\mathcal{D})=\sum_{k=1}^{K}p_{M,T}(\tilde{y}|\tilde{x},\mathcal{D},I_{k})p_{M,T}(I_{k})$
We extend this by introducing the concept of an internal model that the LLM is updating and then using to make its predictions in-context, parameterised by $\varphi$ for each task description:
$p_{M,T}(y|\tilde{x},\mathcal{D})\\ =\int_{\Phi}\sum_{I\in\mathcal{I}}p_{M,T}(y|\tilde{x},\mathcal{D},\varphi)p_{M,T}(\varphi|\mathcal{D},I)p_{M,T}(I)d\varphi$
Where $\varphi$ is itself described by a prior distribution $p_{M,T}(\varphi|I)$, a posterior distribution $p_{M,T}(\varphi|\mathcal{D},I)$, and is unseen. This $\varphi$ is unseen because when we perform in-context learning on a language model for a task description $I$, we only receive $\tilde{y}\sim p_{M,T}(y|\tilde{x},\mathcal{D},I)$ in return.
The inference used to sample $\tilde{y}$ can be performed multiple times to build a sampling distribution that approximates the posterior predictive distribution on $\tilde{y}$ given $\tilde{x}$ and $\mathcal{D}$: $p_{M,T}(\tilde{y}|\tilde{x},\mathcal{D})$.
In the above notation, it is assumed that the output from the language model contains only the predicted label $\tilde{y}$. Alternatively, the log-likelihood of the tokens representing the possible labels can be measured from the language model’s output, leading to a modified notation as in Dong et al. [19].
Learning from demonstrations in-context has proven to be effective for several natural language tasks, demonstrating the surprising ability of language models to perform outside of their intended use [18, 19, 25, 26, 15]. In Brown et al. [18], the authors study GPT-3’s strong ability on natural language tasks after seeing a few examples, one example, or none at all. This work inspired others to study the extent to which language models can learn from demonstrations given as text alongside a task description. Further, Xie et al. [27] study the mechanics of in-context learning and describes it from a Bayesian Inference perspective. They find that in-context learning succeeds when language models can infer a shared concept across individual examples in a prompt. Min et al. [21] argue that in-context learning allows the language model to observe the label space, distribution of input text, and the task format, as they find that swapping the real labels in demonstrations for random labels does not significantly reduce accuracy. The majority of the existing works focus on natural language tasks [19], leaving numerical problems less explored and a focus of our work.
In Hollmann et al. [28], the authors train a transformer model to solve synthetically generated classification tasks from a prior over tabular datasets. Given a new task and demonstrations, this transformer performs in-context learning to make predictions using inference steps only. Additionally, Akyürek et al. [29] and Von Oswald et al. [30] discuss the types of learning algorithms that language models (and more generally, transformers) can implement in-context. Focusing on linear models, Akyürek et al. [29] prove that by construction, transformers are able to perform gradient descent and closed-form ridge regression. Akyürek et al. [29] also demonstrate that trained in-context learners closely match predictors trained through gradient descent. Further, Von Oswald et al. [30] show that transformers, by incorporating multilayer perceptrons into their architectures, can also solve non-linear regression tasks, and Dalal and Misra [31] propose a Bayesian framework to reason about a language model’s internal state and suggest that in-context learning is consistent with Bayesian learning. In our work, we test empirically, through the approximation of the internal model used by a language model for its in-context predictions, whether for a given task, the language model is performing Bayesian inference before and after seeing demonstrations.
To fully compare in-context learning and prior elicitation, we must consider two main questions: (1) Does the language model use the same prior in-context that we elicit? (2) Does in-context learning approximate Bayesian inference? To answer these, we aim to extract, in the case of a linear model, the in-context predictive model’s prior and posterior parameter distributions.
2.3.2 Our contribution: Extracting the in-context prior and posterior
If we assume that the class of the internal predictive model used by the language model can be specified through a prompt (Appendix A.3), then we can estimate the distribution on the parameters of the internal in-context model, $p_{M,T}(\varphi|\mathcal{D})$ and $p_{M,T}(\varphi)$, through the sampling distribution of a maximum likelihood estimator.
By specifying the structure of the internal predictive model to be used by the language model $M$ within our task descriptions $I$ for a task $T$, we can write the internal model as: $f_{\varphi}:X\rightarrow Y$ where $f$ represents the class of the predictive model, $\varphi$ is the random variable associated with its parameters, and $X$ and $Y$ correspond to the predictive model’s domain and range respectively. Depending on the model class $f$ corresponds to, $y\in Y$ could be a probability distribution over a set of possible targets or the prediction of the target itself.
Performing inference on our language model (when it is using its internal predictive model for in-context predictions) returns a $y$ given some $x$. Therefore, although we do not have access to $\varphi$ directly, we do have access to $p_{M,T}(y|x,\varphi)$ and we can perform maximum likelihood estimation to generate a sampling distribution for $p_{M,T}(\varphi)$ and $p_{M,T}(\varphi|\mathcal{D})$.
Given a language model $M$ and task $T$, we construct the sample distribution (following Murphy [32]) as:
$\varphi_{\text{MLE}}\sim\text{MLE}_{M,T}(f_{\varphi}(X)|X,I)$ (2)
Where $\text{MLE}_{M,T}(f_{\varphi}(X)|X,I)$ refers to the maximum likelihood estimate of the parameters $\varphi$ given input feature $X$ and in-context predictions $f_{\varphi}(X)$ for a task description $I$. This is a sample from the estimated prior distribution over the parameters $\varphi$.
We can additionally provide $\mathcal{D}$ examples with the task description to get the posterior samples:
$\varphi_{\text{MLE}}|\mathcal{D}\sim\text{MLE}_{M,T}(f_{\varphi}(X)|X,\mathcal{D},I)$ (3)
These are distributions estimating $p_{M,T}(\varphi)$ and $p_{M,T}(\varphi|\mathcal{D})$. For a more detailed breakdown of the process, see Appendix A.2. In our experiments, to sample each $\varphi_{\text{MLE}}$, we choose to sample $X\sim\mathcal{U}[-5,5]$ five times for each of the $100$ task descriptions. Since we normalise our real data, these samples of $X$ will cover the domain of $f_{\varphi}$ with a high probability.
In the linear case (for both linear and binary logistic regression), this process is simple since we can calculate the maximum likelihood estimate of $y=\varphi_{0}+\varphi\cdot x$ in closed form as we have direct access to $p_{M,T}(y|x,\varphi)$, which is either a regression label or the logit of a class.
Since we now have access to estimates of both $p_{M,T}(\varphi)$ and $p_{M,T}(\varphi|\mathcal{D})$, we can investigate whether the language model is accurately conveying it’s internal prior when it provides us with a prior distribution through elicitation and, when doing in-context learning, whether it is performing Bayesian inference.
2.3.3 Our contribution: Is the language model using the elicited prior in-context?
Since we now have access to an approximation of the prior on a language model’s own internal predictive model $p_{M,T}(\varphi)$, we can study how “truthful” the language model is when it provides us with a prior for a separate model $p_{M,T}(\theta)$ through elicitation (as in Section 2.2 and 3.1). By studying these differences, we can understand one of the trade-offs between using in-context learning and prior elicitation; the language model may not be able to reliably convey its own prior knowledge of a predictive task when we elicit a distribution, or is unable to translate its own prior knowledge into prior distributions over the parameters of its internal predictive model. Either case leads to differences in the predictive models.
To quantify this, we use the energy statistic [33] (Appendix A.4), which measures how closely two distributions match (taking values $\in[0,1]$ with 00 indicating two distributions are equal). Here, we test the difference between the parameter distributions from either the elicited distribution ($\theta\sim p_{M,T}(\theta)$) or the one approximated from the language model’s in-context predictive model ($\varphi\sim p_{M,T}(\varphi)$).
2.3.4 Our contribution: Inspecting the in-context posterior distribution
As part of the decision between using in-context learning and prior elicitation, we would like to understand whether the language model is performing Bayesian inference when it makes predictions in-context.
We can use the approximated internal prior $p_{M,T}(\varphi)$ along with Monte Carlo methods [34, 35] (Appendix A.8) to sample the posterior, and compare it to the approximated internal model posterior $p_{M,T}(\varphi|D)$ presented in Figure 5 and extracted using the MLE sampling introduced in Section 3.3. The difference in these distributions can be measured using the energy statistic as in Section 3.4 to understand whether the language model is acting predictably after observing data from $\mathcal{D}$.
If the language model’s internal posterior is significantly different to the one calculated by applying Monte Carlo methods to the approximated prior, then in-context learning might be considered unusable for a given task, as it is not clear that the model is updating its prior belief about a predictive task from the data. This would suggest that the language model should not be used to generate posterior distributions on predictions in-context as it is improperly applying Bayes rule after observing demonstrations.
2.4 Bayes factor for model selection
To decide, given a dataset to use in-context learning or prior elicitation, we consider the Bayes factor for model selection.
The Bayes factor compares the ratio of marginal likelihoods of two statistical models to determine under which model the data was most likely generated. Formally, given two models $\alpha_{0}$ and $\alpha_{1}$ with parameterisations $\theta_{0}$ and $\theta_{1}$, the Bayes factor calculated over a dataset $\mathcal{D}$ is:
[…]
In our case, $\alpha_{0}$ refers to prior elicitation, with the prior determining the parameterisation $\theta$ of a predictive model, and $\alpha_{1}$ referring to in-context learning. We can approximate the Bayes factor by using the means from samples of the likelihood of $D$ (Section 3.6).
We chose the Bayes factor over the Bayesian information criterion (BIC) because BIC uses the number of model parameters as part of the model selection. When comparing prior elicitation with in-context learning, BIC would almost always favour prior elicitation because the number of parameters required to make a prediction using in-context learning is on the order $10^{9}$.
This method of measurement is only applicable if the language model is performing Bayesian inference when learning in-context, which is explored in Section 3.5.
2.5 Experimental datasets
Within our experiments we present results on five real-world datasets and a synthetic dataset, covering both classification and regression tasks from various domains.
$y=2x_{1}-x_{2}+x_{3}$ is used to generate a synthetic dataset which contains 3 features with values sampled from a normal distribution (mean $=0$ and standard deviation $=1$) and targets calculated using the equation $y=2x_{1}-x_{2}+x_{3}+\epsilon$, where $\epsilon\sim\mathcal{N}(0,0.05)$. When describing this dataset to the language model, we fully specify the relationship between the features and targets to test the setting where the task description is fully descriptive.
To demonstrate the effectiveness of the LLM elicited priors on new datasets, we use a private (i.e., not seen by the LLM) dataset collected as part of a study on Dementia care which contains $10$ features of in-home activity and physiology and labels corresponding to days with a clinically validated labelling of positive or negative urinary tract infection (UTI), calculated through the analysis of a provided urine sample in consultation with a healthcare professional. For this analysis, we used a subset of the data containing $78$ people and $371$ positive or negative days of UTI. This dataset has not appeared online and will not have been part of the language model’s training.
We then test four public benchmark datasets. The Breast Cancer dataset [36], where the task is to predict the diagnosis of a breast mass; the California Housing dataset [37], with the goal to predict the median house value from the U.S census data for that region; the Wine Quality dataset [38, 39] where the task is to predict the wine quality from its physicochemical characteristics; and the Heart Disease dataset [40, 41] where the task is to predict a heart disease diagnosis from physiological features.
Further information on the contents of these datasets is given in Appendix A.5.1. In Appendix A.9, we apply two of the memorisation tests proposed in Bordt et al. [42] to explore the extent to which the language model has memorised the public datasets. The synthetic and UTI datasets are generated or private, and therefore are unseen to the language model.
3 Results
All experiments in this section were conducted with OpenAI’s GPT-3.5-turbo, with further language models tested in Appendix A.13.1 and A.13.2.
3.1 Evaluating the LLM-elicited priors
Figure 2: Test accuracy of the posterior distribution for varied numbers of training data points. Each of the Gaussians in the mixture is extracted from a language model and is tested against an uninformative prior $\mathcal{N}(\theta|0,1)$ on a linear model. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. These values are the result of $5000$ sampled parameters and $10$ random splits of the training and testing data. For California Housing, a regression task, we report the mean squared error on log-scale. 2(a) shows the histogram of accuracy achieved whilst 2(b) shows a line plot of the average mean posterior accuracy or Mean Squared Error (MSE), with the error bars representing the 95% confidence interval, as well as the additional datasets tested in Appendix A.6. The green arrows point in the direction of improvement in the metric.
We start by exploring the performance difference between a predictive model using an uninformative prior distribution over our parameters, $\theta\sim\mathcal{N}(\theta|0,1)$, with one elicited from a language model. For each dataset, we perform cross-validation with five splits and calculate the posterior performance as we vary the number of training data points.
Figure 2(a) shows the results of these experiments, demonstrating that the prior distribution on the linear model’s weights elicited from a language model provides better mean accuracy than an uninformative prior for smaller numbers of data points and often provides less variation in accuracy.
On the synthetic data, we learn to model $y=2x_{1}-x_{2}+x_{3}$ using a prior elicited from the language model. In this case, as we are testing how well the language model can provide priors based on informative task descriptions, we describe the exact relationships between $x$ and $y$ in the task descriptions. The LLM-elicited prior provides a lower MSE, by an order of magnitude, compared to an uninformative prior on the test set after $5$ observed examples and an equivalent MSE after $20$ data points are seen. This is likely because the task was simple enough to be learnt effectively by an uninformative prior after a few examples. However, it demonstrates that the language model is able to provide effective prior distributions given a clear task description. In fact, the mixture of the priors that the language suggests are shown in Appendix A.6.1 and are centred on the true parameter values, suggesting that the language model is able to convert its combined knowledge of a problem and a task description into useful prior distributions. The relationship between the detail of descriptions and the performance of elicited priors is studied within Appendix A.10.
For the real world datasets, we often see the same posterior performance improvements. On the Breast Cancer dataset, the improvement in accuracy is clear up to $40$ observed data points, where the mean accuracy is greater, and its variation is smaller. On the other hand, for the Wine Quality dataset, Figure 2(b) shows that the prior distribution on $\theta$ elicited from the language model is too strict as it eventually reduces the mean accuracy of the logistic regression model when the number of training points is larger than $25$, whilst the variation in accuracy remains smaller. Given these samples, the likelihood of the prior distributions could have been measured (as is described in 2.4) before the elicited distribution was used for new predictions.
Interestingly, for Heart Disease, there is no noticeable difference between the average accuracy of a logistic regression model under an uninformative or elicited prior. This is because the elicited prior contains many Gaussian distributions of mean 00 and standard deviation $1$, making it very similar to the uninformative prior. This can be seen in Figure 11 within Appendix A.7 which plots the elicited parameter distributions as a histogram in each feature dimension, for the first $10$ features. This is reassuring, since it shows that the language model is able to respond with an uninformative prior when it is not confident in its own prior knowledge of a task.
For both the California Housing and UTI datasets, the elicited prior provides a greater mean accuracy over the uninformative prior for all of the number of observed data points tested, illustrating the potential of using a language model for prior elicitation. This should be of particular note considering that the UTI dataset is private, and will have never appeared within the language model’s training.
In general, we see that the greatest improvement in accuracy occurs for smaller numbers of observed samples, suggesting that using prior knowledge from language models is most relevant when users have access to only a few samples.
This can be important in scenarios when collecting labels is costly and time-intensive, as is the case in the UTI dataset. In Appendix A.6.2, we consider the accuracy achieved using uninformative priors and elicited priors in the context of collection time. We show here that a UTI predictive model might have achieved greater accuracy earlier in the study. In a clinical environment, this could have enabled improved UTI detection earlier, accelerating the model development and validation.
In Appendix A.6, we present further accuracy distributions (similar to Figure 2(a)), using the Heart Disease and Wine Quality datasets, which supplement the results in Figure 2(b). We further test the number of task descriptions required in Appendix A.11, and the quality of those descriptions in A.10.
Given two sets of priors, one elicited from a language model and an uninformative prior, the prior predictive distribution can be sampled and the likelihoods calculated given some data points. This can be done similarly to that described in Section 3.6.
3.2 Incorporating expert information
Figure 3: UTI prior parameters with expert information. The histograms, in the three left hand plots, show the distribution of parameter values for the features that we provided expert information for. In all three cases, we stated that the three features are positively correlated with UTI risk. In the right hand plot, we show the posterior accuracy of these prior distributions for different numbers of observed samples. The line shows the mean posterior accuracy and the error bars represent the 95% confidence interval over the $10$ test splits.
By including expert information in the task description, acquired through conversations with domain experts, we can steer a language model to provide prior distributions that combine knowledge from clinicians with its own learnt insights.
To study this, we chose to additionally include our own knowledge of the risk factors for UTI in the task description for prior elicitation [43]. In particular, we provided further information about three features available in the dataset which are positively correlated with outcome. We added the following phrasing to the system prompts, based on clinical knowledge of UTI: “It is known that the more previous UTIs a patient had, the more likely they are to have a UTI in the future. Someone with a UTI will also wake up more during the night. Someone with a UTI will have a greater bathroom usage.”
Figure 3 shows the effect on the prior parameter distribution this has, where all three features we mention in the prompt become more positively predictive of a UTI.
It is interesting however, that without this expert information, the LLM had already correctly identified the relationship between UTI risk and the bathroom frequency, night-time disturbances, and the number of previous UTIs, as these are all positive predictors of UTI. The language model had already translated this information from its internal prior into the elicited distributions. This could explain the small improvement in the posterior accuracy over the LLM-elicited priors without additional expert information, shown in the line graph of Figure 3.
In Appendix A.10, we explore the use of including expert information in the task for a synthetic dataset, where we incrementally include more detailed information.
3.3 Extracting the in-context prior and posterior
Figure 4: Extraction of the language model’s hidden prior for in-context predictions. The bottom row shows the mean squared error between the logits or regression predictions outputted by the language model and the values calculated by the fitted MLE models, whilst the top row shows histograms of the parameter values of these MLE models, with each colour representing a single feature. For ease of visualisation, we exclude values outside of the $(2.5\%,~{}97.5\%)$ percentiles and the bias term. In the top row, each colour represents a single feature.
Figure 5: Extraction of the language model’s hidden posterior for in-context predictions. The bottom row shows the mean squared error between the logits or regression predictions outputted by the language model and the values calculated by the fitted MLE models, whilst the top row shows histograms of the parameter values of these MLE models, with each colour representing a single feature. For ease of visualisation, we exclude values outside of the $(2.5\%,~{}97.5\%)$ percentiles and the bias term. In the top row, each colour represents a single feature.
This section explores the prior and posterior distributions of the linear predictive model that the LLM is using in-context. Further experimental details are available in Appendix A.5.3.
The prior and posterior distributions estimated through a repeated sampling of the in-context predictions (as outlined in Section 2.3.2) are shown in Figure 4 and 5. Once we estimate the parameters of the linear model, we can use them to calculate the logits (or regression values) of the predictions returned by the language model (from in-context learning) and measure the error between the outputs of the language model and the estimated internal model. This is shown in the bottom row of Figure 4 and 5, which suggests that our estimations of the prior internal parameters are reliable for the real-world datasets. The upper row presents the distributions of the estimated parameters, with each colour representing a single feature.
For the prior distributions in Figure 4, we see variations in the estimated prior distributions across the datasets, suggesting that the language model is using some knowledge about tasks when making predictions, even without any demonstration examples. Even so, we see that the language model is often using a prior centred around 00, suggesting it is not particularly informative.
The extracted posterior distributions are given in Figure 5 and show some differences compared to the prior distributions. For California Housing, Heart Disease, and Wine Quality, our estimation error increased, but remained low. However, we find that the Breast Cancer dataset produces a large error for our predictions of the logits (Figure 5), suggesting that our estimates of the parameters of the internal model are not reliable for this case. It is interesting that on the whole, we were able to produce a better approximation of the language model’s internal prior distribution than the posterior distribution, and is likely due to the limitations of the language model and in particular, its capacity to understand demonstrations and instructions in the same prompt.
We do not estimate the language model’s posterior distribution for the UTI dataset as we keep the samples in this dataset private and are therefore unable to provide demonstrations in-context.
3.4 Is the language model being truthful?
Table 1: Extracted prior compared to the elicited prior. The energy statistic between the language model’s internal prior extracted through MLE sampling and the prior elicited by following the method in Section 2.2. Values are $\in[0,1]$ with values closer to zero, indicating that the two distributions are more similar.
UTI Breast Cancer California Housing Wine Quality Heart Disease
0.39 0.31 0.19 0.13 0.18
Table 1 illustrates the differences between the extracted in-context prior distribution (Section 2.3.2) and the LLM-elicited distribution (Section 2.2). No dataset has similar distributions according to their energy distance, demonstrating that the language model is using a different prior distribution to make its own predictions compared to the one it provides for us to construct a separate predictive model.
These conclusions are supported in Appendix A.7, where we present histograms of the parameter values for each feature, which show that the two distributions significantly differ from each other visually, with few similarities. This suggests that either the language model is unable to provide us with a reliable representation of its own prior knowledge or that it is unable to reliably use its own prior knowledge when making predictions. Since the in-context prior distributions are often centred around 00, we suspect that it is the latter, and that the language model is not effectively using its own prior beliefs about a task in-context.
Since these priors disagree, we will investigate them further in Section 3.5, where we explore empirically whether in-context learning is approximating Bayesian inference and in Section 3.6, where we study decision criteria for understanding which of these distributions is more likely to have generated the observed data using the Bayes factor.
3.5 Is in-context learning approximating Bayesian inference?
Table 2: Difference between extracted posterior and MC on extracted prior. The energy statistic between the language model’s internal posterior extracted through MLE sampling and the posterior samples from applying Monte Carlo techniques to the language model’s internal prior. Values are $\in[0,1]$ with values closer to zero, indicating that the two distributions are more similar. Here, we show the mean and standard deviation of the energy statistic for the $5$ different training sets tested as observed data.
Breast Cancer California Housing Wine Quality Heart Disease
0.01 ± 0.00 0.32 ± 0.09 0.02 ± 0.00 0.04 ± 0.01
It is important that the model is performing Bayesian inference, so that newly observed data points and the prior knowledge are combined to form a posterior predictive distribution. In this section, we empirically evaluate the extent to which the language model is performing Bayesian inference in-context, by measuring the difference between the extracted in-context posterior distribution and Monte Carlo sampling applied to the extracted prior in-context distribution (Section 2.3.4).
The results of this analysis are given in Table 2, which shows to what extent the posterior distribution extracted from the language model’s in-context model agrees with Monte Carlo sampling on the extracted in-context prior.
For the Breast Cancer dataset, these two distributions match closely and show a strong agreement in their parameter values, suggesting that for this dataset, in-context learning could be performing Bayesian inference. However, this should be considered alongside the results in Section 3.3, where we found that our approximations of the posterior distribution for the Breast Cancer task had a large error.
For California Housing, we find that we can reliably approximate the in-context prior and posterior models but find a large difference between the sampled posterior (using Monte Carlo methods) and the extracted posterior (using MLE sampling). For this task, the language model failed to perform Bayesian inference in-context, and may suggest the use of prior elicitation instead.
On the other hand, for Wine Quality and Heart Disease, the difference between the extracted posterior and the posterior samples of the extracted prior match closely, and the error in our estimations of the parameters are low (Figures 4 and 5). In these cases, we are more confident that in-context learning is imitating Bayesian inference and so should use a model selection tool to decide between prior elicitation and in-context learning.
Appendix A.8 presents histograms of each feature’s parameter values after Monte Carlo sampling or in-context learning have been performed. Figure 12 within Appendix A.8 supports the findings in this section and allows us to visualise the differences illustrated by Table 2.
These results provide empirical evidence that in-context learning is sometimes approximating Bayesian inference but not necessarily for all tasks.
3.6 Bayes factor and model selection
Figure 6: Distribution of the prior accuracy or mean squared error, the marginal likelihood, and the Bayes factor. The top row shows the distribution of prior accuracy, whilst the bottom row shows the distribution of prior likelihoods. These were calculated using $25$ data points from each dataset, over $5$ test splits and $100$ samples of prompts. The Bayes factor is then calculated between prior elicitation and in-context learning, with a number larger than $1$ favouring prior elicitation. Note that for $y=2x_{1}-x_{2}+x_{3}$ and California Housing (regression tasks), results are plotted on the log-scale, and we exclude values outside of the $(2.5\%,~{}97.5\%)$ percentiles.
To compare prior elicitation with in-context learning (when in-context learning is approximating Bayesian inference), we can calculate the Bayes factor as in Section 2.4.
In the case of prior elicitation, we test the model as described in Section 2.2 and Equation 1. We approximate the log-likelihood of this model by sampling the prior predictive log-likelihood of $\mathcal{D}$. For this, we use the same prior distribution $p_{M,T}(\theta)$ presented in 3.1 and sample $500$ predictions for $25$ data points, repeated with $5$ different splits of the dataset, to get samples of $p(\mathcal{D}|\alpha_{0})$
For in-context learning, we do not have access to $\varphi$ of the internal predictive model. However, since the language model provides us with predictions on targets given features directly, we can still approximate $p(\mathcal{D}|\alpha_{1})$ through a repeated sampling of predictions. Here, we prompt the language model with our $100$ task descriptions and ask it to predict on $25$ data points from our dataset. We repeat this with $5$ different splits of the dataset to calculate the mean and standard deviation of the Bayes factor.
Figure 6 shows the distribution of prior accuracy and marginal likelihood for in-context learning, prior elicitation, and an uninformative prior along with the calculated Bayes Factor.
For the Breast Cancer and Wine Quality datasets, prior elicitation is preferred over in-context learning, whereas for the Heart Disease dataset, in-context learning performs better. In the case of Breast Cancer and Wine Quality, this is likely because the prior distributions elicited from the language model were more informative compared to the one that the language model used in-context. For Heart Disease, the opposite is true, since when performing prior elicitation, the language model returned an uninformative prior (Section 3.1), whilst it used a prior with a smaller parameter range in-context (Section 3.3).
For the regression tasks, prior elicitation is preferred for $y=2x_{1}-x_{2}+x_{3}$, whilst for California Housing in-context learning produces greater log-likelihoods. However, here, both in-context learning and prior elicitation provide a high prior mean squared error and low marginal likelihood, making the Bayes factor unreliable to compute. Further, in the case of California Housing in-context learning is not reliably estimating Bayesian inference (Section 3.5), suggesting prior elicitation may still be preferable.
Whether prior elicitation or in-context learning provides prior distributions with greater likelihoods depends on the task, with both methods being the better choice over a non-informative prior for all cases tested. However, when factoring in the methods’ abilities to estimate Bayesian inference and the cost of producing predictions, prior elicitation is likely the better option for the cases tested.
3.7 Choosing between priors
As seen in Appendix A.13.1 and A.13.2, the prior distributions elicited from language models do not always provide the best posterior performance. We consistently find that GPT-3.5-turbo and GPT-4-turbo provide better prior distributions than an uninformative prior, however this might not be the case for all datasets or language models. Given prior distributions elicited from different language models, we would like to choose the best one for our predictive task. For this, given some collected data, we can use the Bayes factor (Section 2.4) as a model selection technique to decide between the prior distributions elicited from different language models. Alternatively, we could use other model selection techniques such as the Bayesian information criterion. This is because we are performing the Bayesian inference ourselves, given the prior distributions. Further, this method also allows us to compare the elicited distributions to the uninformative prior to ensure the priors from our language models do not harm predictive performance.
By using a temperature parameter, it is also possible to control the strength of the elicited prior, which can indicate our trust in the distribution.
4 Limitations and future research
Within this work, we treat the language model as a domain expert to elicit prior distributions. These prior distributions need to be better than an uninformative prior to improve model performance and reduce data collection costs. A natural next step therefore is to question how these prior distributions compare to those elicited from domain experts, and whether language models provide an additional opportunity to reduce costs by supplementing expert elicited priors or by improving the expert elicitation framework. We began to explore how experts could be asked to produce more detailed task descriptions (Section 3.2 and Appendix A.10), but further settings would be interesting to explore.
As priors elicited from language models are a product of their training, they could contain unwanted bias. An approach to avoid this, would be to assign all features that could be considered sensitive to unwanted bias with an uninformative prior. However, prior distributions on other features might indirectly cause unwanted bias in a predictive model, and it is important to carefully think about whether these features should also be assigned an uninformative prior. Given an elicited prior distribution and observed examples, the Bayes factor can again be used for model selection to choose between combinations of an elicited prior with some or all features replaced with an uninformative prior. Additionally an expert, by providing information in the task description, could steer the language model away from unwanted bias in its provided distributions.
An interesting question is how we extend our results to neural networks. An immediate option is to ask a language model to suggest appropriate data augmentation techniques for a task, effectively reducing the cost of data collection by extending the current dataset. A further exploration of methods would be beneficial.
5 Further results
Within the Appendix, we supplement our findings with further results and discussions. Firstly, we present the full prior elicitation and model selection process on a new dataset in Appendix A.5.5. Then we present additional results of prior elicitation and in-context learning (Appendix A.6, A.7, A.8), where we support the conclusions drawn in Sections 3. Next, since some of the datasets tested are in the public domain and might have appeared in the language model’s training, in Appendix A.9, we provide a discussion on whether the language model has memorised the datasets tested.
To further understand the language model’s behaviour to prompts, we study whether the language model is respecting our in-context model class specifications (Appendix A.3) and test how the detail of the task description affects the posterior performance of the elicited priors (Appendix A.10) and how experts can supplement the process. Further to this, we explore how the number of task descriptions affects the priors elicited (Appendix A.11).
Finally, to contextualise the value of prior elicitation from language models, we report the cost of performing prior elicitation (Appendix A.12) and evaluate the quality of priors elicited from other language models with different price points (Appendix A.13.1) as well as open-source models (Appendix A.13.2).
6 Conclusion
In this work, we presented a method for extracting task information from an LLM to build probabilistic linear models with fewer data points. By asking a language model to play the role of a domain expert, we were able to elicit prior distributions for predictive tasks, potentially reducing the cost of data collection and improving the accuracy of our models when we have access to few labelled examples (Section 3.1). As additionally demonstrated in Appendix A.6.2, these prior distributions can significantly impact the delivery of care in healthcare settings, where collecting labels can be costly and time-consuming, but the need for accurate predictive models is great.
We find that the prior distributions elicited from GPT-3.5-turbo achieve better accuracy or mean squared error with fewer data points than an uninformative prior for a range of tasks (Section 3.1). This performance improvement is most visible when few training examples were available, suggesting its use in environments where label collection is challenging. We also find that incorporating expert information in the task descriptions can enable further improvements in predictive accuracy, and that the language model updates its prior distributions as we would expect (Section 3.2 and Appendix A.10).
Since our approach to prior elicitation can also be seen as a direct replacement for in-context learning for datasets with numerical features, we present a framework for deciding whether, given a task, to use prior elicitation or in-context learning based on the Bayes factor (Section 3.6). However, the Bayes factor assumes that the in-context learner is performing Bayesian inference. We therefore use maximum likelihood estimation and sampling distributions to approximate the internal model being used during in-context learning and its prior and posterior distributions (Section 3.3). This allows us to inspect the truthfulness of the language model when we conduct prior elicitation (Section 3.4), as well as empirically question whether the language model is performing Bayesian inference on a given task (Section 3.5).
We find that prior elicitation is an improved alternative to in-context learning in our setting, as we can inspect priors, confidently perform Bayesian inference, and often achieve reduced error. Prior elicitation also extends the use of the language model’s prior knowledge of predictive tasks to previously infeasible situations, where feature observations might be frequent but labelled examples are few; where in-context learning is too slow or costly.
We have demonstrated that using language models for expert prior elicitation is a cost-effective method for improving the predictive performance of probabilistic linear models in environments with few labels and provides a competitive alternative to in-context learning.
Author Contributions
AC: Original idea conception, software, analysis of results, writing, editing, and reviewing, funding acquisition. RK: Original idea conception, supervision, analysis of results, reviewing. PB: Supervision, analysis of results, reviewing, funding acquisition.
Acknowledgments
This study is funded by the UK Dementia Research Institute (UKDRI) Care Research and Technology Centre funded by the Medical Research Council (MRC), Alzheimer’s Research UK, Alzheimer’s Society (grant number: UKDRI–7002), and the UKRI Engineering and Physical Sciences Research Council (EPSRC) PROTECT Project (grant number: EP/W031892/1). Infrastructure support for this research was provided by the NIHR Imperial Biomedical Research Centre (BRC) and the UKRI Medical Research Council (MRC). Further funding was provided by the Turing scheme. The funders were not involved in the study design, data collection, data analysis, or manuscript writing.
Appendix A Appendix
A.1 Availability of code
Within the supplementary materials we provide code for generating all figures and results, as well as provide a pipeline for using our method on additional datasets.
Our code is implemented in python (v3.11.9) using pip (v24.0).
For our experiments, we use the Python packages:
  • blackjax (v1.2.3) [44]: For Monte Carlo sampling on the in-context prior samples.
  • jax (v0.4.31) [45]: For calculating some metrics and use with blackjax.
  • numpy (v1.26.4) [46]: For manipulation of arrays and computing mathematical equations.
  • openai (v1.51.2): For communicating with the OpenAI API.
  • pandas (v2.2.2) [47]: For manipulating the datasets and results.
  • pymc (v5.17.0) [48]: For Monte Carlo sampling on the elicited priors.
  • scikit-learn (v1.5.1) [49]: For loading and preprocessing the data.
  • transformers (v4.46.0): For loading and using Llama models.
A requirements file is provided within the supplementary code, which also includes some other utility packages required to run the scripts and load the data and results.
A.2 Extracting internal prior and posterior
In this section, we provide a description of the method presented in Section 2.3.2 in more detail.
Given a language model $M$ and task $T$ that provides in-context predictions, if we can construct a function that allows us to sample a single parameterisation $\varphi\sim p_{M,T}(\varphi)$, then we can estimate $p_{M,T}(\varphi)$ using a sampling distribution. Following Murphy [32], we define an estimator:
$\varphi_{\text{MLE}}=\text{MLE}_{M,T}(\hat{\mathcal{D}})\\ =\arg\max_{\tilde{\varphi}}p_{M,T}(\hat{\mathcal{D}}|\tilde{\varphi})$ (5)
Which corresponds to the maximum likelihood estimate (MLE) of $\varphi$ over some dataset $\hat{\mathcal{D}}$ given the task $T$. Now consider that $\hat{\mathcal{D}}$ is generated by a given $\varphi$ sampled from $\varphi\sim p_{M,T}(\varphi)$ and $X$. By sampling $x\sim\mathcal{U}[X]$ from a uniform distribution over $X$, we can create $S$ different datasets of the form:
$\hat{\mathcal{D}}^{(s)}=\{y_{n}\sim p_{M,T}(y_{n}|x_{n},\varphi):n=1:N,~{}x\sim\mathcal{U}[X]\}$ (6)
Which provides us with the sampling distribution:
$\varphi_{\text{MLE}}^{(s)}\sim\text{MLE}_{M,T}(\hat{\mathcal{D}}^{(s)})\\ =\arg\max_{\tilde{\varphi}}p_{M,T}(\hat{\mathcal{D}}^{(s)}|\tilde{\varphi})$
With this, we can generate a sampling distribution using $\{\varphi_{\text{MLE}}^{(s)}\}_{s=0}^{s=S}$ that approximates $p_{M,T}(\varphi)$ as $S\rightarrow\infty$. Additionally, by further incorporating the training dataset $\mathcal{D}$:
$\hat{\mathcal{D}}^{(s)}=\{y_{n}\sim p_{M,T}(y_{n}|x_{n},\varphi,\mathcal{D}):n=1:N,~{}x\sim\mathcal{U}[X]\}$
We get the sampling distribution:
$\varphi_{\text{MLE}}^{(s)}|\mathcal{D}\sim\text{MLE}_{M,T}(\hat{\mathcal{D}}^{(s)}|\mathcal{D})\\ =\arg\max_{\tilde{\varphi}}p_{M,T}(\hat{\mathcal{D}}^{(s)}|\tilde{\varphi},\mathcal{D})$
Allowing us to estimate $p_{M,T}(\varphi|\mathcal{D})$.
A.3 Does the language model use the model class we specify?
Figure 7: In-context predictions for the task: $y=2x_{1}-x_{2}+x_{3}$ compared to the true values. Here, feature values are generated using $X\sim\mathcal{U}(-5,5)$ and the targets are calculated using $y=2x_{1}-x_{2}+x_{3}$, which is specified fully in the in-context prompt. The language model is then asked to make predictions on $X$. This shows the true targets $y$ against those predicted by the language model, as well as the binned root mean squared error.
As part of comparing prior elicitation to in-context learning, we use maximum likelihood sampling to extract the internal parameter distributions (discussed in Section 3.3), which we use to test whether the language model is: (1) providing a similar prior distribution through elicitation to the one it is using in-context, and (2) empirically test whether the language model is approximating Bayesian inference. However, when extracting the internal parameters of the in-context model, we assume that the language model is applying the model class that we specify in the task description.
However, we find that for some datasets, the language model does not use the same model in context as the one we specify. For the synthetic data in particular, we specify the relationship as $y=2x_{1}-x_{2}+x_{3}$. Here, we can measure whether the language model is using the specified model in-context to predict the target by calculating the error between the true $y$ values and those returned by the language model’s in-context predictions. This is done by measuring the mean squared error between the $y_{n}$ given in Equation 6 and the true $\hat{y}_{n}=2x_{n,1}-x_{n,2}+x_{n,3}$ calculated using the $x_{n}$ in Equation 6.
In Figure 7 we show this true $\hat{y}_{n}$ against the in-context prediction $y_{n}$. Interestingly, even given the simple linear equation we ask the language model to mimic, we find that it struggles to provide good predictions in-context. This suggests that the language model is not always using the model class for in-context predictions that we specify in the task description, indicating that in its current state, in-context learning might be inadequate for predictive tasks with primarily numerical features. This figure also shows that the language model was hesitant to predict negative values and preferred positive predictions – a strange behaviour that clearly restricts its use as an emulator of a linear model.
However, for all other prior extraction tasks presented in Section 3.3, we find that the language model uses a linear model in-context, demonstrated by the low values of MSE loss between the predictions of our MLE samples and the true in-context predictions, shown in Figure 11. This is often not the case when the language model is presented with demonstration samples, as in the posterior extraction presented in Figure 5, in which the language model uses a linear combination of features to predict the targets for Heart Disease and Wine Quality only.
A.4 Energy distance and statistic
The energy statistic [33], which is based on the energy distance metric [50], is a method for measuring the similarity of two probability distributions. Importantly, it allows us to test whether two $N$-dimensional probability distributions are equal.
The energy distance $D(X,Y)$, with $X$ and $Y$ as independent random vectors $\in\mathbb{R}^{n}$ is defined as:
$D(X,Y)^{2}=2\mathbb{E}[M_{XY}]-\mathbb{E}[M_{XX}]-\mathbb{E}[M_{YY}]$
Where ${M_{XY}}_{i,j}=||x_{i}-y_{j}||_{2}$ is the pairwise distance matrix.
From this, we construct the energy statistic as:
$H(X,Y)=\frac{D(X,Y)^{2}}{2\mathbb{E}[M_{XY}]}$ (7)
This statistic takes values between 00 and $1$, and tests whether two statistical distributions are equal.
We provide an implementation of this distance metric and test statistic within our supplementary code (Appendix A.1).
A.5 Further experimental details
In all experiments, we use OpenAI’s GPT-3.5-turbo (specifically GPT-3.5-turbo-0125) with a temperature of $0.1$ and all other settings in the OpenAI API default  [^2] https://platform.openai.com/docs/api-reference/chat.
A.5.1 Dataset details
Within this work we evaluate our method on $6$ datasets; $1$ of which is synthetic, $1$ is privately collected, and $4$ are publicly accessible. This allows us to provide both reproducible experiments and test the language model’s ability to suggest a good prior on datasets not seen during training.
  • $y=2x_{1}-x_{2}+x_{3}$: To evaluate language models as tools for prior elicitation, we construct a synthetic dataset where we have control over the parameters of the dataset. Because of this, we are able to provide the language model with a complete description of the task. This dataset contains 3 features sampled from a normal distribution (mean $=0$ and standard deviation $=1$) and targets calculated using the equation $y=2x_{1}-x_{2}+x_{3}+\epsilon$, where $\epsilon\sim\mathcal{N}(0,0.05)$. When training and testing were done on these samples, no preprocessing was performed.
  • A private dataset collected as part of a study on Dementia care: This dataset contains $10$ features engineered from measurements of in-home activity and physiology. In this dataset, the labels corresponded to days with a clinical labelling of positive or negative urinary tract infection (UTI) which were calculated through the analysis of a provided urine sample in consultation with a healthcare professional. For this analysis, we used a subset of the data containing $78$ people and $371$ labelled days of UTI, with $270$ negatives and $101$ positives. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
  • The Breast Cancer dataset [36]: This dataset is made up of $569$ data points with $30$ features computed from an image of a breast mass that are labelled as either malignant or benign. In total, this dataset contains $357$ examples of benign tumours and $212$ examples of malignant tumours. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
  • The California Housing dataset [37]: This dataset contains $20640$ examples of house values in California districts and $8$ features detailing information from the U.S census data for the property’s region. The goal of the dataset is to predict the median house for that region. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
  • The Wine Quality dataset [38, 39]: This dataset contains $6497$ data points of features relating to physicochemical characteristics of red and white wines with the goal of predicting whether the wines are of good quality or not. This dataset contained $2384$ examples of bad quality wine and $4113$ examples of good quality wine. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
  • The Heart Disease dataset [40, 41]: This dataset contains $10$ physiological features and targets with a positive or negative diagnosis of heart disease for $299$ people. Within this dataset, we have $161$ examples of people without heart disease and $138$ examples of people with heart disease, totalling $299$ data points. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data, except where features are categorical. In the categorical case, we perform no scaling.
Where datasets are publicly available, we perform minimal processing for ease of reproducibility. In some cases, we replace the feature names with ones that are more descriptive (taken from the dataset descriptions also available publicly) to aid the language model’s comprehension of the data. All of these changes to the raw data are available within the supplementary code, which will automatically download and process the data upon loading.
A.5.2 Evaluating prior elicitation
To evaluate the usefulness of the priors elicited from the language model (presented in Section 2.2), we first studied their posterior performance in comparison to using an uninformative prior. To do this, we chose the tasks outlined in Section 2.5, $y=2x_{1}-x_{2}+x_{3}$, Breast Cancer, California Housing, Urinary Tract Infection (UTI), Wine Quality, and Heart Disease and calculated the posterior accuracy or mean squared error on a test set after observing varied numbers of examples. For each of the datasets, we selected $50\%$ of data points to form the test set and used the remaining data points to form a training set. This is done $10$ times to produce $10$ folds of random training and testing splits.
For the UTI dataset, we split the dataset such that each participant’s data is only in either the testing or training set, as some participants produced multiple examples. For the other classification tasks, splits are stratified using the targets. To test the posterior performance after observing the given number of data points in the experiments, we randomly select examples from the training set.
After performing prior elicitation (Section 3.1), for each of the tasks, we have $100$ Gaussian priors for each feature in the dataset. We then build a mixture of the priors using Equation 1, where we weigh the Gaussian distribution in each dimension and from each call to the language model separately.
We then construct a logistic or linear regression model, where we set the prior distribution over the bias term as a mixture of $100$ distributions of $\theta_{0}\sim\mathcal{N}(0,1)$. In the case of linear regression, we also use a noise term $\epsilon\sim\text{Half-Cauchy}(\beta=1)$, with a Half-Cauchy prior.
In the uninformative case, we set $\theta\sim\mathcal{N}(0,1)$ for each feature.
We then use the No-U-Turn sampler [35] to sample from the posterior distribution over the linear model parameters, with $5$ chains and $5000$ samples per chain. In total, for each dataset, for each training and testing split, we have $25000$ posterior samples over the posterior.
To produce the results in Figure 2(a) and 8, we then sample the posterior predictive distribution over our test set and calculate the accuracy or mean squared error for each posterior parameter sample. These metric values are then presented in Figure 2(a) and 8, which show their histograms. To produce Figure 2(b) and 13, we also calculate the mean accuracy and mean squared error for each of the test splits before plotting their mean and standard deviation.
For the synthetic task, one of the example task descriptions is as follows:
  • System role: You’re a linear regression predictor, estimating the target based on some input features. The known relationship is: ’target’ = 2 * ’feature 0’ - 1 * ’feature 1’ + 1 * ’feature 2’. Use this to predict the target value.
  • User role: I am a data scientist working with a dataset for a prediction task using feature values to predict a target. I would like to apply your model to predict the target for my samples. My dataset consists of these features: [’feature 0’, ’feature 1’, ’feature 2’]. All the values have been standardized using z-scores. The known relationship is: ’target’ = 2 * ’feature 0’ - 1 * ’feature 1’ + 1 * ’feature 2’. Based on the correlation between each feature and target, whether positive or negative, please guess the mean and standard deviation for a normal distribution prior for each feature. I need this for creating a linear regression model to predict the target. Provide your response as a JSON object with the feature names as keys, each containing a nested dictionary for the mean and standard deviation. A positive mean suggests positive correlation with the outcome, negative for negative correlation, and a smaller standard deviation indicates higher confidence. Only respond with JSON.
A.5.3 Extracting the in-context model’s hidden prior and posterior
To further study in-context learning, we used maximum likelihood sampling to extract approximate samples of the internal model’s prior and posterior using the method described in Section 2.3.2. For each of the tasks studied, we randomly generate $25$ feature values from $X\sim\mathcal{U}[-5,5]$ $5$ times over and ask the language model to predict their regression label or positive classification probability based on the task description given and with the model class specified (in our case a linear or logistic regressor). This is done for each of the $100$ task descriptions we have for the dataset. We then fit a linear regressor to the predicted regression labels or the classification logits  [^3] Calculated using the inverse logistic function: $\text{logits}=\ln p/(1-p)$ of the predicted probabilities, using maximum likelihood estimation. The parameters of this linear model are then taken as a single sample of the approximated in-context predictive model.
For each of these linear models, we then calculate the regression labels or the classification logits using the same data given to the language model for in-context predictions. This allows us to calculate an error in our estimation of the in-context model, which depends on how well the in-context learner is approximating a linear model.
In total, for each task description, we collect $5$ samples of the in-context model parameters. Once this is repeated for all of the $100$ task descriptions, we have $500$ samples of the in-context model.
We can then measure, as in Section 3.4, whether the parameter samples are similar to those sampled from the elicited distributions. This is done by sampling $10000$ parameter values from the elicited distributions and calculating the energy statistic (Appendix A.4) between this sample and the extracted in-context sample. This is done to test whether the language model is using the same distribution we elicit from it, as the one it uses in-context. The results of this are presented in Table 1.
When extracting the posterior distribution over the in-context learner, we repeat the above process $5$ times, where in each repeat we provide the language model with $25$ demonstration points from the corresponding dataset. This allows us to empirically question whether the language model is approximating Bayesian inference when it updates the in-context prior distribution with the observed data points.
To do this, we start by fitting a Gaussian kernel density estimator (with a bandwidth factor of $0.25$) to our in-context model prior parameter samples. We then use the No-U-Turn sampler [35] with $1000$ adaption steps and the same demonstration examples given during in-context learning to sample posterior parameters. Here, we use $100$ chains with $10000$ samples each to build an approximate posterior distribution.
We can then use the energy statistic to measure the difference between the extracted in-context posterior distribution and the Monte Carlo approximation of the posterior based on the extracted in-context prior distribution.
The results presented in Table 2 show the result of this experiment.
A.5.4 Calculating Bayes factor
As part of our solution for deciding whether to use prior elicitation or in-context learning for a given dataset, we calculate the Bayes factor (presented in Section 2.4).
For a single dataset, we start by randomly selecting $25$ data points, $5$ times, for which we consider the prior predictive distribution. For prior elicitation this involves sampling the prior predictive distribution (drawing $500$ samples for each split), and for in-context learning this involves prompting a language model for a prediction for each of the $100$ task descriptions (producing $100$ samples for each of the splits).
For both sets of prior predictive samples, we calculate their log-likelihood using either cross-entropy loss or mean squared error, depending on whether the task is classification or regression. For each dataset, for each test split, for each method, we calculate the mean of this log-likelihood over the samples. With these values, for each dataset and test split, we measure the difference between the value for the elicited prior predictions and the in-context predictions to get the log Bayes factor.
We then report the mean Bayes factor over the splits along with the standard deviation in Figure 6, which helps us to decide, given a dataset, which method is most appropriate.
A.5.5 Given a new dataset
In all of our experiments, we asked the language model to mimic a linear model when predicting in-context, and to provide priors for a linear model when performing prior elicitation.
To apply the prior elicitation techniques to a new dataset, we recommend the following steps:
  1. 1.
    First, create a description of the dataset along with a description of the language model’s role. As part of this, provide descriptive features and target names. It will be helpful to split this task description into what the language model is expected to do (system role) and the task that the user is giving it (user role). Please see our collection of task descriptions in the supplementary code or the example given in Appendix A.5.2.
  2. 2.
    Ask a language model to rephrase your user role and system role many times. In our experiments, we asked the language model to rephrase both descriptions $10$ times for each.
  3. 3.
    Take the product of the system roles and user roles and provide each to the language model. For each of the unique system and user role combinations, record the mean and standard deviation of the Gaussian prior elicited from the language model.
  4. 4.
    Use these individual Gaussian priors to build a mixture of Gaussians as a prior for a linear model.
  5. 5.
    Using the data points you have available from the predictive task, sample from the posterior predictive distribution to predict on unseen data.
Now, to measure whether prior elicitation or in-context learning is better for a given task (assuming the language model is approximating Bayesian inference well):
  1. 1.
    Firstly, sample the prior predictive distribution, with the elicited priors found above, on the data points already collected.
  2. 2.
    Then, to perform in-context learning, use a similar task description to above but with a different system role, that specifies the language model should make predictions instead of providing prior distributions. With this new set of descriptions, again ask a language model to rephrase them.
  3. 3.
    To sample the prior predictive, using the product of these descriptions, ask a language model to make predictions on the data directly.
  4. 4.
    Measure the Bayes factor between the likelihood of the prior predictive distribution given using prior elicitation and the in-context model on the already collected data.
  5. 5.
    Using the Bayes factor, decide whether in-context learning or prior elicitation is more appropriate.
To empirically check whether the language model is approximating Bayesian inference in its in-context predictions or to test whether it is being faithful to the priors we elicit, we can follow the below procedure:
  1. 1.
    Using the same task descriptions as above, provide the language model with randomly generated feature values (we used $25$ samples from $X\sim\mathcal{U}[-5,5]$) for each task description and ask it to make predictions in-context.
  2. 2.
    By additionally providing the seen data as demonstrations to the language model, again ask the language model to make predictions on randomly generated feature values for each task description.
  3. 3.
    With the randomly generated feature values and the in-context predictions, use maximum likelihood estimation to estimate the underlying linear model for each task description. For the classification case, this is more easily done if the language model is asked to provide the probability of a positive prediction rather than the prediction itself.
  4. 4.
    The resulting linear models approximated from the in-context predictions can now be used to approximate the prior and posterior distribution of the in-context model.
Using the above, we can now (1) empirically test whether the language model is being faithful to the prior distributions we elicit and (2) empirically test whether the language model is approximating Bayesian inference in-context.
For the former, we can use a measure of the difference in samples (in our case, the energy distance, Appendix A.4) to calculate the difference between the in-context model’s prior distribution and the elicited prior distribution. This will inform us of whether the language model is being truthful when it provides us with a prior on the parameters of a linear model, or whether the language model uses that prior knowledge when making predictions in-context. If the prior distribution elicited differs from the one that the language model is using in-context, then calculating the Bayes factor could be important to ensure that we select the most appropriate predictive model for the given task. If it is not clear whether the language model is approximating Bayesian inference in-context, we should then study the latter to ensure the Bayes factor is a suitable model selection tool.
Therefore, for the latter, we can follow the below procedure:
  1. 1.
    Use a kernel density estimator (KDE) to approximate a smoother version of the prior distribution over the in-context model. For our experiments, we used a Gaussian kernel density estimator with a bandwidth factor of $0.25$.
  2. 2.
    Using the KDE and Monte Carlo methods, given the collected data, we sample a posterior distribution from the in-context learner’s prior.
  3. 3.
    Use a method to calculate the the difference between two samples of a distribution to understand the difference between the posterior samples calculated through Monte Carlo sampling of the in-context prior and the samples of the in-context posterior.
This last procedure might indicate that our in-context learner is not approximating Bayesian inference and is not suitable for a given task, where it could be important to us that our predictive model is updating its prior knowledge in an intuitive way.
A.6 Further prior elicitation results
Figure 8: Test accuracy of the posterior distribution over the model weights of a logistic and linear regression model for varied numbers of training data points. We evaluate the accuracy and mean squared error of different prior distributions over the weights of a logistic and linear regression model. Each of the Gaussians in the mixture is extracted from a language model and is tested against an uninformative prior $\mathcal{N}(\theta|0,1)$. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. These values are the result of $100$ sampled priors and $5$ random splits of the training and testing data. Note that the bottom row shows the mean squared error (on log-scale) where lower values are better. The green arrows point in the direction of improvement in the metric.
Here, we present the results of the same experiment as discussed in Section 3.1, applied to two further datasets discussed in Section 2.5.
The results of a logistic regression model on a test set after observing training data and using a prior collected through elicitation of the language model or an uninformative prior are shown in Figure 8.
These results suggest that for the Wine Quality dataset, the prior elicited is better than the uninformative prior for smaller numbers of observations but is too strict when the number of observations is larger since the result distribution is more rigid than for an uninformative prior. This is similar to what is observed in the results on the Breast Cancer and UTI datasets presented in Section 3.1.
Further, for the Heart Disease dataset, we find that the prior elicited from the language model performed comparably to the uninformative prior. This is because the language model provided mean = 0 and std = 1 for a large number of priors elicited, suggesting that the language model was unsure about the prior distribution on the parameters for a model predicting heart disease. The elicited parameter distributions can be seen in Figure 11, which contextualises these results.
A.6.1 Priors elicited from the synthetic data
Figure 9 shows a histogram of parameter values from the GPT-3.5-turbo elicited prior for the task descriptions of $y=2x_{1}-x_{2}+x_{3}$. Here, the language model correctly provides prior distributions that are centred around the correct parameter values and offer a small standard deviation.
Figure 9: Elicited priors for the task $y=2x_{1}-x_{2}+x_{3}$. We sample $10000$ parameter values from the mixture generated from the Gaussian priors elicited from the language model when given the task description. This figure shows the histograms of those values in each feature dimension.
A.6.2 UTI accuracy with the label collection context
Figure 10: Posterior accuracy of the elicited priors on the UTI dataset considering label collection time. The average mean posterior accuracy and 95% confidence interval on the UTI test set after observing the given number of data points against the time it took for the data points to be collected in the study.
Since we have access to the underlying study that collected the UTI dataset, we are able to understand in more detail how the accuracy of the predictive model varies in time, as well as with the number of data points collected.
In Figure 10 we plot the length of time it took to collect the given number of labels tested in the experiments in Section 3.1 against the accuracy of a model using an uninformative prior and one elicited from a language model. This result illustrates the usefulness of prior elicitation, since we are able to achieve comparable (and ultimately better) accuracy with less data points, and therefore less data collection time.
Since UTI label collection is costly and slow, our UTI predictive model could have achieved greater accuracy significantly earlier in the study. This is paramount in a clinical environment, where an improved UTI detection model could have resulted in fewer unplanned hospital admissions.
It is important to note, that the accuracy in Figure 10 is not achieved as a result of training on the earliest collected data points. This is because in our analysis, we perform cross-validation and use a subset of the data from later time period (Section A.5.1). There were changes to the in-home devices during the study, such as the addition of a device for collecting sleep information, which we use within the dataset evaluated in the main results. However, if all of the data collection devices had been used from the beginning of the study, we might expect the accuracy in time to resemble the figure presented here.
A.7 Is the language model being truthful when providing priors through elicitation?
Figure 11: Distribution of the prior parameters under prior elicitation and those extracted from the internal model for the different predictive tasks. Here, we compare the prior distributions of the different predictive tasks, as defined by the internal model used in-context, and that which we elicit from the language model. Here, we draw the histogram separately in each dimension, but calculate the energy distance across all features combined.
In Section 3.4, we study how truthful the language model is when we elicit a prior distribution, compared to the prior it uses in its internal predictive model when performing in-context learning.
Figure 11 visualises the different prior parameter distributions for the UTI, Breast Cancer, California Housing, Wine Quality, and Heart disease datasets in each feature dimension. Of note is the large differences between the elicited prior distributions and the ones used when performing in-context learning, which supports the findings in Section 3.4. In the in-context case, the prior distributions frequently have their modes close to 00, suggesting the internal model is unsure about the relationship between features and targets. On the other hand, the elicited priors often represent a stronger belief about the correlation between features and targets, with most of the histograms achieving modes greater than or less than 00.
Interestingly, on the Breast Cancer dataset, prior elicitation resulted in two distinct modes in the prior distribution for Feature 2 (“mean texture”), which could represent uncertainty in the direction of correlation.
Further, the prior distribution on the parameters for the Heart Disease dataset have more variation in their values in comparison to that extracted from the language model’s internal model. This tells us that the elicited prior parameters represent more uncertainty in their value and are therefore more flexible when calculating the posterior on observed data.
A.8 Testing empirically whether the language model is approximating Bayesian inference
In this section, we visually inspect the differences between the posterior distribution extracted through maximum likelihood sampling of the language model’s internal model (Section 3.3) and one that is sampled through MCMC methods on the extracted prior distribution.
Figure 12: Distribution of the posterior parameters extracted from the internal model for the predictive tasks against the posterior MC samples with extracted prior for a single data split. Here, we compare the distribution of the posterior extracted from the language model against the distribution that we sample using MCMC methods on the extracted prior distribution. Both posterior distributions are based on the same $25$ data point sample from the corresponding dataset. We draw the histogram separately in each dimension, but calculate the energy distance across all features combined. This is calculated using one of the $5$ training sets tested, hence the difference in values compared to Table 2.
For a single random draw of $25$ training data points from each corresponding dataset, in Figure 12, we sample parameter values from the posterior distributions and present them as a histogram in each feature dimension.
In the cases of Breast Cancer, Wine Quality, and Heart Disease, we see that the posterior distributions overlap significantly and show a strong agreement, which is confounded by the small value of the energy statistic between the histograms calculated across all features combined. This is empirical evidence that the language model is performing Bayesian inference in-context, since it is aligned with the posterior estimates from an MCMC method.
On the California Housing dataset, the posteriors do not match as closely. The histograms given in Figure 12, along with the comparatively larger energy distance, show a significant difference in the distributions. In the extracted posterior, almost all of the values are contained within a small range around zero, whilst the posterior sampled from MCMC methods shows a larger variation in possible values. Additionally, in this cases, the energy statistic is far greater than for other datasets and suggests that for the California Housing regression task, in-context learning is not approximating Bayesian inference.
A.9 Dataset memorisation
To better contextualise the use of in-context learning on the datasets tested, we will utilise the method proposed in Bordt et al. [42] to quantify the extent to which the language model already knows information about the dataset.
Table 3 shows the results of the header completion test described in Bordt et al. [42], which tests whether the language model has memorised the first few rows of the dataset. The column names, along with a given number of the first rows are given to the language model as a string, with the final of these rows truncated. The language model is then asked to generate the next $500$ tokens.
We use the Levenshtein distance [51, 52], a string matching metric, to calculate the agreement between the true header completion and the one provided by the language model. The Levenshtein distance counts the number of single-character changes required to transform one string into another, only allowing for insertions, deletions, and substitutions. The value can take a minimum value of 00 – suggesting the strings are equal and a maximum value of the greatest length of the two strings.
To make the Levenshtein distance comparable across datasets, we calculate its normalised value by dividing the Levenshtein distance by the length of either the true string or the language model’s completion, whichever is longer. Because of this, the normalised Levenshtein distance reported is not formally a metric as it does not follow the triangle inequality. However, it allows us to compare the header completions across datasets more easily:
$\hat{L}(s_{1},s_{2})=\frac{\text{Levenshtein}(s_{1},s_{2})}{\text{min}\{|s_{1}|,|s_{2}|\}}$
Where $\text{Levenshtein}(\cdot,\cdot)$ calculates the Levenshtein distance between two strings, $\text{min}\{\cdot,\cdot\}$ returns the minimum value of the inputs, and $|s|$ refers to the length of $s$. This new value ranges between 00 and $1$, with 00 indicating a perfect match between the strings and $1$ meaning that they were different at all positions.
Table 3: Header completion test results. Normalised Levenshtein distance between the correct header completion and the language model’s completion for each of the datasets tested. A value of 00 indicates that the true completion and language model’s completion were equal and a value of $1$ means that they were different at all positions.
Breast Cancer California Housing Wine Quality Heart Disease
0.00 0.15 0.21 0.00
The values in Table 3 demonstrate that the language model has memorised the first few rows of the Breast Cancer and Heart Disease datasets completely, demonstrated by a normalised Levenshtein distance value of 0. This could be because the first few rows of these datasets commonly appear within the language model’s training data. For California Housing and Wine Quality, we see some level of memorisation, with the language model being able to reproduce a large amount of the true header, however, to a much lesser extent than what we see for Breast Cancer and Heart Disease. As we might expect, when we perform the header test on the synthetic dataset generated using the relationship $y=2x_{1}-x_{2}+x_{3}$, we get a Normalised Levenshtein distance of $0.7$, which shows a large difference between the true header string and the language models completion. The UTI dataset has not been tested, as it contains sensitive information and is currently a private dataset.
Table 4: Row completion test results. Normalised Levenshtein distance between the correct row completion and the language model’s completion for each of the datasets tested. A value of 00 indicates that the true completion and language model’s completion were equal and a value of $1$ means that they were different at all positions.
Breast Cancer California Housing Wine Quality Heart Disease
0.43 ± 0.13 0.58 ± 0.04 0.31 ± 0.04 0.33 ± 0.05
We supplement these results with those in Table 4, which show the results of the row completion test. Here we calculate the normalised Levenshtein distance of the true completions and the language model’s completion of a random row from the dataset. In these tests, the language model is provided with $10$ rows (in order from the original dataset) and is asked to complete the following row. The values in Table 4 represent the mean and standard deviation of the test using $25$ random rows. Here, in contrast to the header test, we see that the language model has not memorised all rows from the dataset. This is because the language model’s completion differs from the true row completion in all cases to varying degrees.
This is likely because when the datasets appear within the language model’s training data, their column names and initial rows are given, but not the entire dataset.
A.9.1 How does this affect our results?
Within our experiments we normalise the training and testing data, as well as sample random observations from the dataset. Because of this, observations predicted during in-context learning will be likely not have been memorised, since they are normalised based on a random training data subset, and are selected from random rows of the dataset. As shown, although the language model is able to complete the first few rows of the datasets with some success, it is unable to complete random rows (Table 4). Therefore, we hypothesise that any memorisation of the tested datasets has little impact on the in-context predictions.
Further, since for prior elicitation we only prompt the language model for a prior distribution of a predictive model given the feature names and a description of the dataset, the memorisation of rows from the data is unlikely to have had an impact on the results. In addition, we see that elicited priors provide improved predictions over the uninformative prior for the private or synthetic datasets, as well as the public datasets.
A.10 Task descriptions with more information
Figure 13: Mean squared error using the elicited priors for the task: $y=2x_{1}-x_{2}+x_{3}$ where we have used increasingly more detailed task descriptions. Here, feature values are generated using $X\sim\mathcal{U}(-5,5)$ and the targets are calculated using $y=2x_{1}-x_{2}+x_{3}$. Within each task description we provide the language model with increasingly more information about the task, as to test how the task description effects the priors elicited. Complete descriptions of the task descriptions given here are provided in Section A.10. The lines depict the average mean posterior mean squared error (MSE) over the 5 dataset splits, whilst the error bars represent the 95% confidence interval.
To understand how the task description affects the performance of the priors elicited from the language model, we studied the synthetic task, $y=2x_{1}-x_{2}+x_{3}$, and described the relationship between features and targets in varied detail. We additionally test an uninformative prior to contextualise the results. For each of the levels of detail shown in Figure 13, we included the following text in the task description that was given to the language model:
  • Linear relationship:
    “The target is linear in features”
  • One feature relationship:
    “The target is a linear combination of the features and that when ‘feature 0’ increases by 1, the target increases by 2”
  • Two feature relationships:
    “The target is a linear combination of the features and that when ‘feature 0’ increases by 1, the target increases by 2, and when ‘feature 1’ increases by 1, the target decreases by 1”
  • Three feature relationships:
    “The target is a linear combination of the features and that when ‘feature 0’ increases by 1, the target increases by 2, and when ‘feature 1’ increases by 1, the target decreases by 1, and when ‘feature 2’ increases by 1, the target increases by 1”
  • Full equation:
    “The ‘target’ = 2 * ‘feature 0’ - 1 * ‘feature 1’ + 1 * ‘feature 2’”
For each prior elicited using these levels of detail, in Figure 13, the corresponding $\log_{10}\text{MSE}$ on the test set can be seen for a linear regressor after observing the given number of data points. This figure illustrates the effects of varied descriptions and shows that the improvements made to the MSE diminish as the task description contains more detail. The biggest improvement came from providing the language model with a single feature relationship – enabling an MSE that is two orders of magnitude better.
Providing the language model with improved levels of detail allows it to produce prior distributions that enable our linear regression model to achieve a given error rate with fewer data points. As we might expect, the more detail provided in the task description, the better our prior distributions are.
What is unexpected, however, is that for this task, using an uninformative prior ($\theta\sim\mathcal{N}(\theta|0,1)$) provides better posterior performance than specifying to the language model that the relationship between the target and features is linear (Figure 13). This is explained by inspecting the elicited prior distribution when being given this information, which shows that the language model responded with a similar mean and standard deviation for a Gaussian for each of the $100$ variations of prompts:
  • Feature 0: mean: 0.5 ± 0.11, std: 0.12 ± 0.05
  • Feature 1: mean: -0.32 ± 0.08, std: 0.08 ± 0.06
  • Feature 2: mean: 0.56 ± 0.26, std: 0.14 ± 0.09
This demonstrates the sometimes unexpected behaviour of language models, which here produced Gaussians for the prior distribution with small standard deviations (reflecting a high confidence in the distribution), and means that imply the language model knew the underlying equation defining the relationship between the target and features. We might have expected that after only being told the relationship was linear, the language model would produce Gaussians with 00 mean and standard deviation $1$, reflecting the uncertainty in the exact relationship between each feature and the target.
Interestingly, although “Three feature relationships” and “Full equation” contain the same information written differently, the language model provided better prior distributions when given the full equation. Fully specifying the equation is not realistic in a real-world setting where we are not sure about the exact relationship between features and targets, but this provides an example of how we can make the task description as information-dense as possible.
Figure 14: Parameter distribution as further feature information is included. These histograms show the distribution of parameter values for the features as more expert information is provided. In each row, we provide information in the task description of an additional feature, with the green arrow indicating the effect this has on the parameter distribution. In particular, in the first row we provide information about the relationship between the target and the first feature, in the second row we provide information about the second feature, and the in the third row we provide information about the third feature.
An expert interacting with the language model might be able to provide further information about a task, that improves the elicited priors. In Figure 14 we show the parameter distributions elicited from the language model for the different descriptions we provide. This figure shows that as we provide additional information about each feature, the language model responds by providing a prior distribution that reflects its updated knowledge of the task. This demonstrates the potential of using an expert to construct more detailed task descriptions, which the language model can translate into parameter distributions and supplements the results in Section 3.2.
A.11 How many ways should we describe a task?
Figure 15: Measuring how similar the prior distributions are with fewer task descriptions. We reduce the number of task descriptions and calculate the energy statistic between $10000$ samples from the prior elicited from the language model using the reduced set of descriptions and the full $100$ task descriptions. We repeat each experiment $10$ times and show the mean and standard deviation.
In this section we look at how many of the task descriptions are required to be able to get a similar prior distribution to that elicited within our experiments,using $100$ task descriptions.
Figure 15 shows the energy statistic (Appendix A.4) of samples of the prior distribution elicited using fewer task descriptions compared to samples from the prior distribution elicited using $100$ task descriptions (from the same language model). In particular, we compare a subset of the task descriptions to the full $100$ task descriptions.
We see that for most tasks, a large number of task descriptions is required, since reducing the number of descriptions from $100$ to $49$ produced prior distributions that are significantly different for all datasets tested. Using $81$ task descriptions produces prior distributions that are more similar but still show differences for some datasets (for example $y=2x_{1}-x_{2}+x_{3}$ and Breast Cancer).
It is interesting to see that Breast Cancer benefitted most from increasing the number of task descriptions, since this dataset contains the largest number of features. It suggests that more task descriptions are required for those datasets with high-dimensional data. This makes intuitive sense, since we would expect to require a larger mixture of Gaussian components to better describe the prior over for a higher dimensional dataset.
This figure shows that as the number of task descriptions increases, we see smaller changes in the energy distance, and reduced variability in the samples. This suggests that increasing the number of task descriptions provides diminishing returns.
Further work could allow the number of task descriptions to be a function of the number of features in the dataset.
A.12 Cost of elicitation
The priors elicited from GPT-3.5-turbo incur a cost, as we use OpenAI’s API. To elicit all of the priors used in the experiments in Section 3.1, we were charged \$$\$0.41$ (on average \$$\$0.068$ per dataset). However, we did not make use of the batch API and instead made each API call sequentially. All of the experiments in this work could be implemented using the batch API since no new API call depends on a previous API result, which may reduce cost significantly  [^4] https://openai.com/api/pricing/.
A.13 Other language models
In this section, we will compare the results presented in the main paper (on GPT-3.5-turbo) to some other commonly used language models. We start by exploring the performance of some other OpenAI models  [^5] https://platform.openai.com/docs/models, before considering some of the open source Llama models  [^6] https://huggingface.co/meta-llama [15, 16, 17].
A.13.1 Other OpenAI models
Figure 16: Test accuracy of the posterior distribution for different OpenAI language models. Here we show the results when using priors elicited from two other OpenAI language models and an uninformative prior $\mathcal{N}(\theta|0,1)$ on a linear model. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. These values are the result of $5000$ sampled parameters and $10$ random splits of the training and testing data. For California Housing, a regression task, we report the mean squared error on log-scale. This figure shows the line plot of the average mean posterior accuracy or mean squared error (MSE), with the error bars representing the 95% confidence interval. The green arrows point in the direction of improvement in the metric.
In Figure 16 we also present the posterior results of priors elicited from both GPT-4-turbo and GPT-4o-mini. GPT-4-turbo incurs a significantly higher cost than the GPT-3.5-turbo model used for the main results in this paper, whilst GPT-4o-mini provides a less expensive option.
Specifically, to elicit all priors presented in Figure 16, GPT-4o-mini cost \$$\$0.32$ (on average \$$\$0.053$ per dataset), whilst GPT-4-turbo cost \$$\$7.56$ (on average \$$\$1.26$ per dataset). This is compared to the cost of \$$\$0.41$ (on average \$$\$0.068$ per dataset) for GPT-3.5-turbo (presented in A.12). The more features a dataset contains, the more expensive priors are to elicit since the number of input and output tokens is greater.
Figure 16 demonstrates that in most cases, GPT-4-turbo provides priors that are significantly better than GPT-3.5-turbo but falls behind in the regression tasks. On Breast Cancer, UTI, and Wine Quality, the priors elicited from GPT-4-turbo allow for a significantly better posterior accuracy, and unlike GPT-3.5-turbo, provide an informative prior for Heart Disease. However, for both the regression tasks (the synthetic task and the California Housing task), the GPT-3.5-turbo priors allow for a better mean squared error (MSE) than GPT-4-turbo.
GPT-4o-mini often produces prior distributions that are equal to or worse than GPT-3.5-turbo. For the synthetic task, the priors elicited from GPT-4o-mini are as good as an uninformative prior; for Breast Cancer and UTI, they are as good as GPT-3.5-turbo; for Wine Quality, Heart Disease, and California Housing however, they are worse than GPT-3.5-turbo. For GPT-4o-mini, eliciting priors for Heart Disease caused significant problems, as they were worse than both the uninformative prior and those elicited from GPT-3.5-turbo.
A.13.2 Open source language models
Figure 17: Test accuracy of the posterior distribution for Llama language models. Here, we show the results when using priors elicited from two of the Llama language models and an uninformative prior $\mathcal{N}(\theta|0,1)$ on a linear model. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. These values are the result of $5000$ sampled parameters and $10$ random splits of the training and testing data. For California Housing, a regression task, we report the mean squared error on log-scale. This figure shows the line plot of the average mean posterior accuracy or mean squared error (MSE), with the error bars representing the 95% confidence interval. The green arrows point in the direction of improvement in the metric.
Here, we explore the posterior performance of prior distributions elicited from three Llama [15, 16, 17] models and one Qwen model [53, 54]:
In contrast to the OpenAI models in Appendix A.13.1, these models are publicly available and open-source after being granted access. This means they can be used without the need for connecting to an online API, and when locally hosted, all data and inference can be performed on device, enabling the use of our framework within a closed and secure environment.
The experiments presented in this section use the same task descriptions as given to GPT-3.5-turbo, except that because the Llama models often fail to return a JSON structure, we suffix the prompt with an additional instruction to return a JSON. When sampling the posterior with the elicited priors, we use the same methods and options as given in A.5.2.
Figure 17 illustrates the inconsistent posterior accuracy of the priors elicited from the Llama family of language models tested. These models are significantly smaller in capacity to GPT-3.5-turbo, justifying their inconsistent elicited priors, but in some of the datasets, they still outperform the uninformative prior.
Specifically for Breast Cancer, the 3B parameter Llama model provides an intermediate accuracy between an uninformative prior and those elicited from GPT-3.5-turbo, and for the synthetic task $y=2x_{1}-x_{2}+x_{3}$ and California Housing, the 8B parameter Llama provides priors of equal quality to GPT-3.5-turbo. On the other hand, for the UTI, Wine Quality, and Heart Disease datasets, the Llama models both provide prior distributions that are of worse quality than the uninformative prior. It is possibly not surprising that the Llama-3.2-3B-Instruct provided priors that were equal or worse than Llama-3.1-8B-Instruct in all cases except for predicting Breast Cancer since these models differ significantly in size.
Llama-3.1-70B-Instruct-BF16 is more competitive, sometimes providing priors that allow for better posterior distributions than GPT-3.5-turbo, but on other occasions, providing priors with similar posterior performance to the uninformative prior. On the UTI and Heart Disease datasets, the prior distributions elicited from Llama-3.1-70B-Instruct-BF16 provide better posterior accuracy than those from GPT-3.5-turbo. Surprisingly, on Heart Disease prediction in particular, priors from the 70B Llama model provided a posterior accuracy greater than GPT-4-turbo. Additionally, when predicting Breast Cancer, priors form Llama-3.1-70B-Instruct-BF16 achieved the same posterior accuracy as GPT-3.5-turbo for all of the number of training examples tested. These three datasets suggest that Llama-3.1-70B-Instruct-BF16 can be a useful model for eliciting prior distributions for a predictive task. However, when eliciting priors for the Wine Quality, California Housing, and $y=2x_{1}-x_{2}+x_{3}$ tasks, we found that Llama-3.1-70B-Instruct-BF16 provided distributions that were not competitive with GPT-3.5-turbo.
Qwen-2.5-14B-Instruct is similarly inconsistent in the posterior performance that the priors provide. For Heart disease, it achieves a greater posterior accuracy than all models tested, except for the 70B Llama model, Llama-3.1-70B-Instruct-BF16. For the UTI task, after $10$ demonstrations are seen, it performs comparably to Llama-3.1-70B-Instruct-BF16, and better than GPT-3.5-turbo. On Wine Quality it achieves the greatest posterior accuracy after $20$ observed data points. However, on the California Housing, Breast Cancer, and the synthetic task, it performs worse than GPT-3.5-turbo, but is likely smaller in capacity. Overall, Qwen-2.5-14B-Instruct does not provide consistent performance across the tasks, but has the benefit of performing similarly to Llama-3.1-70B-Instruct-BF16 on many tasks with a significantly smaller compute requirement.
It is important, therefore, given an elicited prior distribution, to calculate the prior likelihood with some already collected data to understand whether that distribution is helpful for constructing a predictive model.