### v3

# Using Large Language Models for Expert Prior Elicitation in Predictive Modelling

Abstract
Large language models (LLMs), trained on diverse data effectively acquire a breadth of information across various domains. However, their computational complexity, cost, and lack of transparency hinder their direct application for specialised tasks. In fields such as clinical research, acquiring expert annotations or prior knowledge about predictive models is often costly and time-consuming. This study proposes the use of LLMs to elicit expert prior distributions for predictive models. This approach also provides an alternative to in-context learning, where language models are tasked with making predictions directly. In this work, we compare LLM-elicited and uninformative priors, evaluate whether LLMs truthfully generate parameter distributions, and propose a model selection strategy for in-context learning and prior elicitation. Our findings show that LLM-elicited prior parameter distributions significantly reduce predictive error compared to uninformative priors in low-data settings. Applied to clinical problems, this translates to fewer required biological samples, lowering cost and resources. Prior elicitation also consistently outperforms and proves more reliable than in-context learning at a lower cost, making it a preferred alternative in our setting. We demonstrate the utility of this method across various use cases, including clinical applications. For infection prediction, using LLM-elicited priors reduced the number of required labels to achieve the same accuracy as an uninformative prior by 55%, 200 days earlier in the study.
1 Introduction
In fields such as clinical research, labelling data can be costly and time-consuming [1]. For example, in healthcare monitoring, patient data is recorded through wearables or remote monitoring devices, and observations are continuously collected. However, labelling such data for clinical outcomes, such as diagnosing illnesses or collecting and studying samples, can be challenging. In addition, estimating the uncertainty in clinical predictions is paramount [2], especially when available training data is sparse.
Expert knowledge of a problem domain can provide prior distributions of model parameters. This can improve a model’s performance when trained with few data points and enable estimates of uncertainties on parameters and predictions. However, expert knowledge about model parameter distributions is often difficult to obtain due to the challenges of performing prior elicitation from experts [3, 4], especially in low-resource settings.
Due to their diverse training data, language models offer an opportunity to begin resolving these issues. By considering a language model as an “expert” in a given field, we can extract insights about a predictive task that might lead to improved predictive models. This, and the abilities of language models to reason about problem settings outside of their intended use cases [5, 6, 7, 8, 9, 10], have inspired their use in performing prior elicitation [11, 12, 13]. Specifically, Selby et al. [12] show, by considering prior predictive distributions, that GPT-4 [10] can provide parameter distributions that allow for zero-shot performance (making predictions without training examples) of Bayesian models that compare to models trained on tens of data points. Furthermore, in Requeima et al. [13], the authors prompt language models with text and collected data points to elicit posterior predictive distributions for regression tasks, which they find outperforms Gaussian processes with radial basis function kernels on some tasks. This is done by repeatedly sampling a language model at a new feature value, given the previous input and output observations, and a given text description that contains prior information from a human expert. Given the rate of development of language models [14, 15, 16, 17], it is conceivable that their knowledge improves and broadens with time. This motivates the construction of a probabilistic framework to extract prior task knowledge from language models and evaluate the use of their elicited distributions for predictive tasks.
Figure 1: Prior elicitation using a language model.
In-context learning allows for the direct use of language models for predictive tasks. Here, language models are provided with task descriptions and demonstration examples from a training data set (in text form) and are asked to make predictions on unseen data [18, 19]. In this sense, in-context learning is similar to eliciting a prior distribution on a predictive model’s parameters and calculating its posterior predictive distribution, where the main differences arise from: (1) Language models are notoriously weak at performing arithmetic on numerical values [20]; (2) Language models lack transparency and may not be learning from demonstrations in a predictable or Bayesian way [21]; (3) Language models may not be able to faithfully provide their internal knowledge of a task when specifying a prior parameterisation for a separate predictive model or use that same prior knowledge in-context. (4) Language models incur a significant cost, and so their use for predictions in a high-frequency or low-resource setting is infeasible. To study the trade-offs between using in-context learning and prior elicitation, we aim to use the Bayes factor for model selection, and propose an approach to estimate a language model’s internal predictive prior and posterior models.
We are motivated by the following questions and make the following contributions:
  1. 1.
    Question: Can language models provide good prior distributions over the parameters of a predictive model, which enables the training of more robust models with less data?
    Method: We propose a method to elicit general prior parameter distributions for linear models from language models for predictive tasks (Section 2.2).
    Finding: We find that using language models for expert prior elicitation provides better posterior accuracy and mean square error than an uninformative prior on a variety of numerical predictive tasks, enabling the construction of predictive models with fewer labelled examples (Section 3.1).
  2. 2.
    Question: How do the LLM’s in-context prior distributions compare to the ones we elicit?
    Method: We present an approach for extracting the in-context prior distribution (Section 2.3.2).
    Finding: We find that these distributions differ significantly from those elicited (Section 3.4).
  3. 3.
    Question: How can we check if the language model is performing Bayesian inference in-context?
    Method: By additionally providing the language model with training demonstrations, we extract the posterior distribution of the LLM’s in-context parameters (Section 2.3.2). We can then perform Monte Carlo sampling on the extracted prior and measure its difference to the extracted posterior.
    Finding: We find that the language model is inconsistent in its approximation of Bayesian inference and that its reliability depends on the task (Section 3.5).
  4. 4.
    Question: When is performance better with in-context learning?
    Method: We use the Bayes factor as a model selection method to decide between prior elicitation and in-context learning given a task (Section 2.4).
    Finding: Before considering cost and reliability, the Bayes factor depends on the task. However, after considering the in-context model’s inconsistent approximation of Bayesian inference and its cost, we conclude that prior elicitation is the preferred choice for all tasks tested (Section 3.6).
To implement our methods and reproduce our results, we provide details on the code [^1] https://github.com/alexcapstick/llm-elicited-priors in Appendix A.1 and complete experimental information in Appendix A.5.
2 Methods
2.1 Defining the task descriptions
The goal of this section is to define the space of prompts that we use to elicit priors and extract in-context predictions from a language model. We start by defining the space of all tasks as $\mathcal{T}$. An individual task, such as a specific classification or regression problem, will be referred to as $T\in\mathcal{T}$. Since a description of a task can be rephrased many times without changing its meaning, we must consider the space of possible task descriptions, referred to as $\mathcal{I}$. We say that the probability of a random text sequence $I\in\mathcal{I}$ being a relevant task description for task $T\in\mathcal{T}$ is: $p(I|T)$.
We assume that all relevant task descriptions have equal probability, that is, if $I_{i}$ and $I_{j}$ fully describe $T$, then $p(I_{i}|T)=p(I_{j}|T)$. We also assume that irrelevant task descriptions have $p(I|T)=0$. This means that for a task $T$, all of the mass of $p(\mathcal{I}|T)$ is uniformly spread across the relevant task descriptions only.
In reality, the number of relevant task descriptions would be too large to cover, since each word in the task description could be replaced with a synonym without losing the meaning of the full description. In Appendix A.11, we show that there are diminishing returns as you increase the number of ways of describing a task. For our experiments, we use $100$ relevant task descriptions to cover this space.
In practice, to produce $100$ task descriptions, we employ a language model to rephrase a single human written system role and user role $10$ times each and take their product. This means that using $100$ task descriptions is not particularly time-consuming and could be significantly increased by providing many base descriptions or prompting the language model to produce many more variants.
2.2 Using language models to elicit expert priors
2.2.1 Background
Bayesian inference allows us to incorporate our knowledge about a predictive task before observing the data. This is done using the prior knowledge $p(\theta)$ (over a predictive model parameterised by $\theta$), which can be our own intuition about which features are more important for a prediction and in what way they correlate. Additionally, using a prior allows us to report uncertainty over new predictions $(\tilde{x},\tilde{y})$, through the posterior predictive distribution $p(y|\tilde{x},D)$. If the prior is well-specified, using a prior can also reduce the number of data points required to construct a comparably accurate predictive model.
A key problem in machine learning is to define this prior distribution $p(\theta)$ given expert knowledge. Previous work studied how this distribution can be elicited from experts [3], which can be challenging and costly, or infeasible in low-resource settings. In parallel, research has shown that large language models (LLMs) encode knowledge about a surprising variety of domains [5, 6, 7, 8, 9, 10], which could be used to generate information about a predictive task [11, 12, 13], possibly supplemented with scientific discussions between experts and the language model[22].
Choi et al. [11] explores the use of an LLM to perform feature selection, causal discovery, and to modify the reward function in a reinforcement learning task. In the first case, they observe that when listing features from two datasets, GPT-3 can correctly identify the features that correspond to a single specified predictive task. Additionally, when applied to causal discovery, the language model is often able to predict the directionality of causal relationships in the Tüebingen Cause-Effect Pairs dataset [23]. This inspires us to explore the possibility of using LLMs for eliciting prior parameter distributions over linear models.
Further, Selby et al. [12] elicit parameters of a beta distribution to define a prior for predictive tasks where results from human experts are available. The priors from four LLMs were visually compared to those elicited from human experts using the SHELF [3] method, which demonstrated significant differences between experts and LLMs. In our work, we extend this by using the LLM to elicit multiple Gaussian distributions as priors. We combine these to create a mixture model for a more general distribution over our parameters and test their utility for linear predictive models.
Furthermore, Zhu and Griffiths [24] describe the use of iterated learning to elicit prior distributions from language models for causal learning, proportion estimation, and predicting everyday quantities. The authors employ a Markov chain Monte Carlo method to access an LLM’s prior knowledge by performing successive inference using previous model outputs that tend towards a limiting distribution. In doing this, the authors assume that in-context learning is a form of Bayesian inference. However, in their work Zhu and Griffiths [24] do not consider how prior elicitation from LLMs can be used for predictive models or test whether an LLM is performing Bayesian inference, both of which we study in this work.
2.2.2 Our contribution: LLM-elicited priors
Our first contribution, therefore, focuses on proposing a method for eliciting an LLM’s prior knowledge of a task to construct general distributions on the parameters of a predictive model that improve performance with less data compared to a predictive model with an uninformative prior and produce uncertainty about its predictions. An overview of this process is given in Figure 1.
Formally, given a language model $M$ and task $T$, we will obtain a single Gaussian prior for each feature, for each task description $I$. This is done by sampling $(\mu_{k},\sigma_{k})$ from $p_{M,T}(\mu,\sigma|I_{k})$, where we obtain a single $\mu_{k}$ and $\sigma_{k}$ by prompting the language model to provide a guess of the mean and standard deviation of a Gaussian prior for each feature on a linear model.
By taking a mixture of these individual Gaussian distributions, we construct a prior $p(\theta|M,T)=p_{M,T}(\theta)$ over our model parameters $\theta$:
$p_{M,T}(\theta)=\sum_{k=1}^{K}\pi_{k}\mathcal{N}(\theta|\mu_{k},{\sigma_{k}}^{2})\\ ~{}~{}\text{where}~{}(\mu_{k},\sigma_{k})\sim p_{M,T}(\mu,\sigma|I_{k})~{}\text{and}~{}\pi_{k}\sim\text{Dir}(1)$ (1)
In our experiments, our mixture has over $100$ components, using several different task descriptions. Mixtures of Gaussians are able to approximate any probability density with enough components, allowing us to approximate the language model’s belief about the task, even if non-Gaussian. We would also expect the language model to have learnt how to parameterise a Gaussian distribution from its training, making them good candidates. However, mixtures of other distributions can be used instead and may be more useful for particular tasks.
The posterior predictive distribution on unseen data $(\tilde{x},\tilde{y})$, given an observed dataset $\mathcal{D}$, becomes:
$p_{M,T}(y|\tilde{x},\mathcal{D})=\int_{\Theta}\sum_{k=1}^{K}p_{M,T}(y|\tilde{x},\theta)p_{M,T}(\theta|\mathcal{D},I_{k})p_{M,T}(I_{k})d\theta$
Here, $p_{M,T}(I)$ refers to the probability distribution discussed in Section 2.1. The distribution $p_{M,T}(\theta|\mathcal{D},I)$ describes the posterior distribution over the parameterisation of a predictive model, which is calculated using the available data $\mathcal{D}$ and the prior parameterisation we extract from the language model.
The practical sampling of these distributions is discussed in Section 3.1 and Appendix A.5.2.
Since we are using a language model’s prior knowledge and a given training dataset to make predictions, this method is an alternative to in-context learning. The difference is that prior elicitation uses a marginal distribution and Bayesian inference to address some of the shortfalls of in-context learning.
2.3 In-context learning
2.3.1 Background
In-context learning aims to use the knowledge that language models acquire during training to make predictions on unseen data. Specifically, this method requires providing a language model with a description of the predictive task and some demonstration examples that allow it to learn “in-context” before the user queries the predictions on some unseen data.
Probabilistically, given a language model $M$ and task $T$, and many task descriptions $I_{k}\in\mathcal{I}$, a prediction of $\tilde{y}$ on features $\tilde{x}$ using in-context learning with observed data $\mathcal{D}$ is:
$p_{M,T}(y|\tilde{x},\mathcal{D})=\sum_{k=1}^{K}p_{M,T}(\tilde{y}|\tilde{x},\mathcal{D},I_{k})p_{M,T}(I_{k})$
We extend this by introducing the concept of an internal model that the LLM is updating and then using to make its predictions in-context, parameterised by $\varphi$ for each task description:
$p_{M,T}(y|\tilde{x},\mathcal{D})\\ =\int_{\Phi}\sum_{I\in\mathcal{I}}p_{M,T}(y|\tilde{x},\mathcal{D},\varphi)p_{M,T}(\varphi|\mathcal{D},I)p_{M,T}(I)d\varphi$
Where $\varphi$ is itself described by a prior distribution $p_{M,T}(\varphi|I)$, a posterior distribution $p_{M,T}(\varphi|\mathcal{D},I)$, and is unseen. This $\varphi$ is unseen because when we perform in-context learning on a language model for a task description $I$, we only receive $\tilde{y}\sim p_{M,T}(y|\tilde{x},\mathcal{D},I)$ in return.
The inference used to sample $\tilde{y}$ can be performed multiple times to build a sampling distribution that approximates the posterior predictive distribution on $\tilde{y}$ given $\tilde{x}$ and $\mathcal{D}$: $p_{M,T}(\tilde{y}|\tilde{x},\mathcal{D})$.
In the above notation, it is assumed that the output of the language model contains only the predicted label $\tilde{y}$. Alternatively, the log-likelihood of the tokens representing the possible labels can be measured from the language model’s output, leading to a modified notation as in Dong et al. [19].
Learning from demonstrations in-context has proven to be effective for several natural language tasks, demonstrating the surprising ability of language models to perform outside of their intended use [18, 19, 25, 26, 15]. In Brown et al. [18], the authors study GPT-3’s strong ability on natural language tasks after seeing a few examples, one example, or none at all. This work inspired others to study the extent to which language models can learn from demonstrations given as text alongside a task description. Further, Xie et al. [27] study the mechanics of in-context learning and describes it from a Bayesian Inference perspective. They find that in-context learning succeeds when language models can infer a shared concept across individual examples in a prompt. Min et al. [21] argue that in-context learning allows the language model to observe the label space, the distribution of input text, and the task format, as they find that swapping the real labels in demonstrations for random labels does not significantly reduce accuracy. Most of the existing work focuses on natural language tasks [19], leaving numerical problems less explored and a focus of our work.
In Hollmann et al. [28], the authors train a transformer model to solve synthetically generated classification tasks from a prior over tabular datasets. Given a new task and demonstrations, this transformer performs in-context learning to make predictions using inference steps only. Furthermore, Akyürek et al. [29] and Von Oswald et al. [30] discuss the types of learning algorithms that language models (and more generally, transformers) can implement in-context. Focusing on linear models, Akyürek et al. [29] show that, by construction, transformers are capable of performing gradient descent and closed-form ridge regression. Akyürek et al. [29] also demonstrate that trained in-context learners closely match predictors trained through gradient descent. Further, Von Oswald et al. [30] show that transformers, by incorporating multilayer perceptrons into their architectures, can also solve non-linear regression tasks, and Dalal and Misra [31] propose a Bayesian framework to reason about a language model’s internal state and suggest that in-context learning is consistent with Bayesian learning. In our work, we test empirically, through the approximation of the internal model used by a language model for its in-context predictions, whether for a given task, the language model is performing Bayesian inference before and after seeing demonstrations.
To fully compare in-context learning and prior elicitation, we must consider two main questions: (1) Does the language model use the same prior in-context that we elicit? (2) Does in-context learning approximate Bayesian inference? To answer these questions, we aim to extract, in the case of a linear model, the in-context predictive model’s prior and posterior parameter distributions.
2.3.2 Our contribution: Extracting the in-context prior and posterior
If we assume that the class of the internal predictive model used by the language model can be specified through a prompt (Appendix A.3), then we can estimate the distribution on the parameters of the internal in-context model, $p_{M,T}(\varphi|\mathcal{D})$ and $p_{M,T}(\varphi)$, through the sampling distribution of a maximum likelihood estimator.
By specifying the structure of the internal predictive model to be used by the language model $M$ within our task descriptions $I$ for a task $T$, we can write the internal model as: $f_{\varphi}:X\rightarrow Y$ where $f$ represents the class of the predictive model, $\varphi$ is the random variable associated with its parameters, and $X$ and $Y$ correspond to the domain and range of the predictive model, respectively. Depending on the model class $f$ corresponds to, $y\in Y$ could be a probability distribution over a set of possible targets or the prediction of the target itself.
Performing inference on our language model (when it is using its internal predictive model for in-context predictions) returns a $y$ given some $x$. Therefore, although we do not have access to $\varphi$ directly, we do have access to $p_{M,T}(y|x,\varphi)$ and we can perform maximum likelihood estimation to generate a sampling distribution for $p_{M,T}(\varphi)$ and $p_{M,T}(\varphi|\mathcal{D})$.
Given a language model $M$ and task $T$, we construct the sample distribution (following Murphy [32]) as follows:
$\varphi_{\text{MLE}}\sim\text{MLE}_{M,T}(f_{\varphi}(X)|X,I)$ (2)
Where $\text{MLE}_{M,T}(f_{\varphi}(X)|X,I)$ refers to the maximum likelihood estimate of the parameters $\varphi$ given input feature $X$ and in-context predictions $f_{\varphi}(X)$ for a task description $I$. This is a sample of the estimated prior distribution over the parameters $\varphi$.
We can also provide $\mathcal{D}$ examples with the task description to get posterior samples:
$\varphi_{\text{MLE}}|\mathcal{D}\sim\text{MLE}_{M,T}(f_{\varphi}(X)|X,\mathcal{D},I)$ (3)
These are distributions that estimate $p_{M,T}(\varphi)$ and $p_{M,T}(\varphi|\mathcal{D})$. For a more detailed breakdown of the process, see Appendix A.2. In our experiments, to sample each $\varphi_{\text{MLE}}$, we choose to sample $X\sim\mathcal{U}[-5,5]$ five times for each of the $100$ task descriptions. Since we normalise our real data, these samples of $X$ will cover the domain of $f_{\varphi}$ with a high probability.
In the linear case (for both linear and binary logistic regression), this process is simple since we can calculate the maximum likelihood estimate of $y=\varphi_{0}+\varphi\cdot x$ in closed form since we have direct access to $p_{M,T}(y|x,\varphi)$, which is either a regression label or the logit of a class.
Since we now have access to estimates of both $p_{M,T}(\varphi)$ and $p_{M,T}(\varphi|\mathcal{D})$, we can investigate whether the language model is accurately conveying it’s internal prior when it provides us with a prior distribution through elicitation and, when doing in-context learning, whether it is performing Bayesian inference.
2.3.3 Our contribution: Is the language model using the elicited prior in-context?
Since we now have access to an approximation of the prior on a language model’s own internal predictive model $p_{M,T}(\varphi)$, we can study how “truthful” the language model is when it provides us with a prior for a separate model $p_{M,T}(\theta)$ through elicitation (as in Section 2.2 and 3.1). By studying these differences, we can understand one of the trade-offs between using in-context learning and prior elicitation; the language model may not be able to reliably convey its own prior knowledge of a predictive task when we elicit a distribution, or is unable to translate its own prior knowledge into prior distributions over the parameters of its internal predictive model. Either case leads to differences in the predictive models.
To quantify this, we use the energy statistic [33] (Appendix A.4), which measures how closely two distributions match (taking values $\in[0,1]$ with 00 indicating that two distributions are equal). Here, we test the difference between the parameter distributions from either the elicited distribution ($\theta\sim p_{M,T}(\theta)$) or the one approximated from the language model’s in-context predictive model ($\varphi\sim p_{M,T}(\varphi)$).
2.3.4 Our contribution: Inspecting the in-context posterior distribution
As part of the decision between using in-context learning and prior elicitation, we would like to understand whether the language model is performing Bayesian inference when it makes predictions in-context.
We can use the approximated internal prior $p_{M,T}(\varphi)$ along with Monte Carlo methods [34, 35] (Appendix A.8) to sample the posterior, and compare it to the approximated internal model posterior $p_{M,T}(\varphi|D)$ presented in Figure 5 and extracted using the MLE sampling introduced in Section 3.3. The difference in these distributions can be measured using the energy statistic as in Section 3.4 to understand whether the language model is acting predictably after observing data from $\mathcal{D}$.
If the language model’s internal posterior is significantly different from the one calculated by applying Monte Carlo methods to the approximated prior, then in-context learning might be considered unusable for a given task, as it is not clear that the model is updating its prior belief about a predictive task from the data. This would suggest that the language model should not be used to generate posterior distributions on predictions in-context as it is improperly applying Bayes rule after observing demonstrations.
2.4 Bayes factor for model selection
To decide whether to use in-context learning or prior elicitation, given a data set, we consider the Bayes factor for model selection.
The Bayes factor compares the marginal likelihood ratio of two statistical models to determine under which model the data was most likely generated. Formally, given two models $\alpha_{0}$ and $\alpha_{1}$ with parameterisations $\theta_{0}$ and $\theta_{1}$, the Bayes factor calculated over a dataset $\mathcal{D}$ is:
[…]
In our case, $\alpha_{0}$ refers to prior elicitation, with the prior determining the parameterisation $\theta$ of a predictive model, and $\alpha_{1}$ referring to in-context learning. We can approximate the Bayes factor using the means from samples of the likelihood of $D$ (Section 3.6).
We chose the Bayes factor over the Bayesian Information Criterion (BIC) because BIC uses the number of model parameters as part of the model selection. When comparing prior elicitation with in-context learning, BIC would almost always favour prior elicitation because the number of parameters required to make a prediction using in-context learning is of the order $10^{9}$.
This method of measurement is only applicable if the language model is performing Bayesian inference when learning in-context, which is explored in Section 3.5.
2.5 Experimental datasets
Within our experiments, we present results on five real-world datasets and a synthetic dataset, covering both classification and regression tasks from various domains.
The equation $y=2x_{1}-x_{2}+x_{3}$ is used to generate a synthetic dataset that contains 3 features sampled from a normal distribution (mean $=0$ and standard deviation $=1$) and targets calculated using the equation and perturbed by noise $\epsilon\sim\mathcal{N}(0,0.05)$. When describing this dataset to the language model, we fully specify the relationship between the features and targets to test the setting where the task description is fully descriptive.
To demonstrate the effectiveness of the LLM elicited priors on new tasks, we use a private dataset (i.e., not seen by the LLM) collected as part of a study on Dementia care which contains $10$ features of in-home activity and physiology and labels corresponding to days with a clinically validated labelling of positive or negative urinary tract infection (UTI), calculated through the analysis of a provided urine sample in consultation with a healthcare professional. For this analysis, we used a subset of data containing $78$ people and $371$ positive or negative days of UTI. This dataset has not appeared online and will not have been part of the language model’s training.
We then test four public benchmark datasets. The Heart Disease dataset [36, 37] where the task is to predict a heart disease diagnosis from physiological features; the Diabetes dataset [38], with the goal of predicting a quantitative measure of disease progression one year after baseline; the Hypothyroid dataset [39, 40], where the task is to predict whether a patient has hypothyroidism based on biomarkers; and the Breast Cancer dataset [41], where the task is to predict the diagnosis of a breast mass.
Further information on the contents of these datasets is given in Appendix A.5.1. In Appendix A.9, we apply two of the memorisation tests proposed in Bordt et al. [42] to explore the extent to which the language model has memorised the public datasets. The synthetic and UTI datasets are generated or private, and therefore are unseen to the language model.
3 Results
All experiments in this section were conducted with OpenAI’s GPT-3.5-turbo, with further language models tested in Appendices A.13.1 and A.13.2.
3.1 Evaluating the LLM-elicited priors
Figure 2: Test accuracy of the posterior distribution for varied numbers of training data points. Each of the Gaussians in the mixture is extracted from a language model and is tested against an uninformative prior $\mathcal{N}(\theta|0,1)$ on a linear model. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. These values are the result of $5000$ sampled parameters and $10$ random splits of the training and testing data. For Diabetes, a regression task, we report the mean squared error on log-scale. 2(a) shows the histogram of accuracy achieved whilst 2(b) shows a line plot of the average mean posterior accuracy or Mean Squared Error (MSE), with the error bars representing the 95% confidence interval, as well as the additional datasets tested in Appendix A.6. The green arrows point in the direction of improvement in the metric.
We start by exploring the performance difference between a predictive model using an uninformative prior distribution over our parameters, $\theta\sim\mathcal{N}(\theta|0,1)$, with one elicited from a language model. For each dataset, we perform cross-validation with five splits and calculate the posterior performance as we vary the number of training data points.
Figure 2(a) shows the results of these experiments, demonstrating that the prior distribution on the linear model’s weights elicited from a language model provides better mean accuracy than an uninformative prior for smaller numbers of data points and often provides less variation in accuracy.
On the synthetic data, we learn to model $y=2x_{1}-x_{2}+x_{3}$ using a prior elicited from the language model. In this case, since we are testing how well the language model can provide priors based on informative task descriptions, we describe the exact relationships between $x$ and $y$ in the task descriptions. The LLM-elicited prior provides a lower MSE, by an order of magnitude, compared to an uninformative prior on the test set after $5$ observed examples and an equivalent MSE after $20$ data points are seen. This is likely because the task was simple enough to be learnt effectively by an uninformative prior after a few examples. However, it demonstrates that the language model is capable of providing effective prior distributions given a clear task description. In fact, the mixture of priors that the language suggests are shown in Appendix A.6.1 and are centred on the true parameter values, suggesting that the language model is able to convert its combined knowledge of a problem and a task description into useful prior distributions. The relationship between the details of the descriptions and the performance of elicited priors is studied within Appendix A.10.
Figure 2(b) shows that for real-world datasets, we often see the same posterior performance improvements. On the Breast Cancer dataset, the improvement in accuracy is clear up to $40$ observed data points, where the mean accuracy is greater, and its variation is smaller. The elicited prior distribution achieved a greatly increased accuracy for lower numbers of observed data points, enabling more reliable predictions earlier into model development.
In the case of Hypothyroid prediction, using the elicited priors did not improve the posterior predictions for low numbers of observed data points. However, once the number of observed data points increased to $20$ and greater, the accuracy was higher when using the elicited priors. As in Section 2.4 and 3.6, with a Bayesian approach, the likelihood of the prior distributions can be measured given the observed data before they are used for new predictions and the appropriate prior can be selected without access to a test set.
Interestingly, for Heart Disease, there is no noticeable difference between the average accuracy of a logistic regression model under an uninformative or elicited prior. This is because the elicited prior contains many Gaussian distributions of mean 00 and standard deviation $1$, making it very similar to the uninformative prior. This is illustrated in Figure 12 within Appendix A.7 which plots the elicited parameter distributions as a histogram in each feature dimension, for the first $10$ features. This is reassuring, since it shows that the language model is able to respond with an uninformative prior when it is not confident in its own prior knowledge of a task.
For both the Diabetes and UTI datasets, the elicited prior provides a greater mean accuracy or lower MSE over the uninformative prior for all of the number of observed data points tested. This should be of particular note considering that the UTI dataset is private and will have never appeared within the language model’s training.
In general, we see that the greatest improvement in accuracy occurs for smaller numbers of observed samples, suggesting that the use of prior knowledge from language models is most relevant when users have access to only a few samples. This can be important in scenarios where collecting labels is costly and time-intensive, as is the case in the UTI dataset. In Appendix A.6.2, we consider the accuracy achieved using uninformative priors and elicited priors in the context of collection time. We show here that a UTI predictive model might have achieved greater accuracy earlier in the study. In a clinical environment, this could have enabled improved UTI detection earlier, accelerating model development and validation.
In Appendix A.6, we present additional accuracy distributions (similar to Figure 2(a)), using the Heart Disease and Hypothyroid datasets, which supplement the results in Figure 2(b). In Appendix A.11, we examine the number of task descriptions required to elicit consistent prior distributions, and how the detail of these descriptions affects posterior performance in A.10.
3.2 Incorporating expert information
Figure 3: UTI prior parameters with expert information. The histograms, in the three left hand plots, show the distribution of parameter values for the features that we provided expert information for. In all three cases, we stated that the three features are positively correlated with UTI risk. In the right hand plot, we show the posterior accuracy of these prior distributions for different numbers of observed samples. The line shows the mean posterior accuracy and the error bars represent the 95% confidence interval over the $10$ test splits.
By including expert information in the task description, acquired through conversations with domain experts, we can steer a language model to provide prior distributions that combine knowledge from clinicians with its own learnt insights.
To study this, we chose to also include our own knowledge of the risk factors for UTI in the task description for prior elicitation [43]. In particular, we provide further information on three features available in the dataset that are positively correlated with the outcome. We added the following phrasing to the system prompts, based on clinical knowledge of UTIs: “It is known that the more previous UTIs a patient had, the more likely they are to have a UTI in the future. Someone with a UTI will also wake up more during the night. Someone with a UTI will have a greater bathroom usage.”
Figure 3 shows the effect on the prior parameter distribution this has, where all three features we mention in the prompt become more positively predictive of a UTI.
It should be noted that without this expert information, the LLM had already correctly identified the relationship between the risk of UTI and bathroom usage, the frequency of nighttime disturbances, and the number of previous UTIs, as these are all positive predictors of UTI. The language model had already translated this information from its internal prior into the elicited distributions. This could explain the small improvement in posterior accuracy over LLM-elicited priors without additional expert information, shown in the line graph of Figure 3.
In Appendix A.10, we examine the elicited priors when incrementally including more detailed expert information in the task description for a synthetic dataset, which a language model holds no prior knowledge for.
3.3 Extracting the in-context prior and posterior
Figure 4: Extraction of the language model’s hidden prior for in-context predictions. The bottom row shows the mean squared error between the logits or regression predictions outputted by the language model and the values calculated by the fitted MLE models, whilst the top row shows histograms of the parameter values of these MLE models, with each colour representing a single feature. For ease of visualisation, we exclude values outside of the $(2.5\%,~{}97.5\%)$ percentiles and the bias term. In the top row, each colour represents a single feature.
Figure 5: Extraction of the language model’s hidden posterior for in-context predictions. The bottom row shows the mean squared error between the logits or regression predictions outputted by the language model and the values calculated by the fitted MLE models, whilst the top row shows histograms of the parameter values of these MLE models, with each colour representing a single feature. For ease of visualisation, we exclude values outside of the $(2.5\%,~{}97.5\%)$ percentiles and the bias term. In the top row, each colour represents a single feature.
This section explores the prior and posterior distributions of the linear predictive model that the LLM is using in-context (Appendix A.5.3).
The prior and posterior distributions estimated through a repeated sampling of the in-context predictions (Section 2.3.2) are shown in Figures 4 and 5. To quantify our approximation error, once we estimate the parameters of the linear model, we can use them to calculate the predictions returned by the language model in-context to measure the difference between the output of the language model and the estimated internal model. This is shown in the bottom row of Figures 4 and 5, suggesting that our estimates of the internal parameters are reliable for the real-world datasets.
For the prior distributions in Figure 4, we see variations in the estimated prior distributions across the datasets, suggesting that the language model uses some knowledge about the tasks when making predictions, even without any demonstration examples. However, we see that the language model often uses a prior centred close to 00, suggesting that it is not particularly informative.
The extracted posterior distributions are presented in Figure 5 and show some differences compared to the prior distributions. For the Diabetes and Heart Disease tasks our estimation error increased but remained low. However, we find that the Breast Cancer and Hypothyroid dataset produces a large error for our predictions of the logits (Figure 5), suggesting that our estimates of the internal posterior model parameters are not reliable for this case. The error in approximating the parameters of the in-context model for the synthetic task is discussed in Appendix A.3. It is interesting that, on the whole, we were able to produce a better approximation of the language model’s internal prior distribution than the posterior distribution. This is likely due to the limitations of the language model, and in particular, its capacity to understand demonstrations and instructions in the same prompt.
As LLMs improve their ability to follow instructions, these approximation errors will reduce considerably since they are a product of the language model’s inconsistent imitation of a linear model.
We do not estimate the language model’s posterior distribution for the UTI dataset as samples in this dataset are kept private, and so we are unable to provide demonstrations in-context.
3.4 Is the language model being truthful?
Table 1: Extracted prior compared to the elicited prior. The energy statistic between the language model’s internal prior extracted through MLE sampling and the prior elicited by following the method in Section 2.2. Values are $\in[0,1]$ with values closer to zero, indicating that the two distributions are more similar.
UTI Heart Disease Hypothyroid Diabetes Breast Cancer
0.39 0.18 0.15 0.38 0.31
Table 1 illustrates the differences between the extracted in-context prior distribution (Section 2.3.2) and the LLM-elicited distribution (Section 2.2). No dataset has similar distributions according to their energy distance, demonstrating that the language model uses a different prior distribution to make its own predictions, to the one it provides for us to construct a separate predictive model.
These conclusions are supported in Appendix A.7, where we present histograms of the parameter values for each feature, which show that the two distributions visually differ significantly. This suggests that either the language model is unable to provide us with a reliable representation of its own prior knowledge or that it is unable to reliably use its own prior knowledge when making predictions. Since the in-context prior distributions are often centred around 00, we suspect that it is the latter and that the language model does not effectively use its own prior beliefs about a task in-context.
Since these priors disagree, we will investigate them further in Section 3.5, where we explore empirically whether in-context learning is approximating Bayesian inference and in Section 3.6, where we study decision criteria to understand which of these distributions is more likely to have generated the observed data using the Bayes factor.
3.5 Is in-context learning approximating Bayesian inference?
It is important that the model performs Bayesian inference, so that newly observed data points and prior knowledge are combined to form a posterior predictive distribution. In this section, we empirically evaluate the extent to which the language model is performing Bayesian inference in-context. This is done by measuring the difference between the extracted in-context posterior distribution and Monte Carlo (MC) posterior samples of the extracted prior in-context distribution (Section 2.3.4).
Figure 6: The different posterior samples Here, we compare the distribution of the posterior extracted from the language model’s in-context predictions against the distribution that we sample using Monte Carlo methods on the extracted prior distribution. Both posterior distributions are based on the same $25$ data point sample from the corresponding dataset. We draw the histogram separately in each dimension, but calculate the energy distance across all features combined. For the diabetes dataset we have removed the long tails for improved visualisation.
The results of this analysis are given in Figure 6, which shows the extent to which the posterior distribution extracted from the language model’s in-context model agrees with Monte Carlo sampling on the extracted in-context prior. In this figure, we present the results for those datasets which achieve a low approximation error when estimating the posterior in-context parameters (Figure 5). The results of the other datasets are given in Appendix A.8, alongside those on other dataset splits.
The results for the two datasets given in Figure 6 present contradictory conclusions. For the Diabetes task, the posterior parameter samples from the in-context posterior and MC sampling on the in-context prior visually agree and produce a low energy statistic, suggesting that the LLM was able to approximate Bayesian inference in-context. On the other hand, for the Heart Disease task, we find that the two methods of sampling posterior parameters yielded different results, implying that for this task, the LLM was not correctly approximating Bayesian inference.
These results provide empirical evidence that whether in-context learning approximates Bayesian inference depends on the task. The LLM inconsistently applies the Bayes rule in context, which suggests that it should not be considered a reliable classifier. In this case, our prior elicitation framework presents itself as the more dependable option.
As the field develops, it is feasible that future LLMs can reliably perform Bayesian inference, and so in these cases we require a model selection tool to decide between different methods for incorporating a language model’s knowledge into a predictive task.
3.6 Bayes factor and model selection
Figure 7: The distribution of the prior marginal likelihood, and the Bayes factor. This shows the distribution of prior likelihoods for the tasks. These were calculated using $25$ data points from each dataset, over $5$ test splits and $100$ samples of prompts. The Bayes factor is then calculated between prior elicitation and in-context learning, with a number larger than $1$ favouring prior elicitation. For the regression tasks, we use the log-scale and exclude values outside of the $(2.5\%,~{}97.5\%)$ percentiles.
To compare prior elicitation with in-context learning, when in-context learning is approximating Bayesian inference, we can calculate the Bayes factor as described in Section 2.4.
In the case of prior elicitation, we test the model as described in Section 2.2 and Equation 1. We approximate the marginal log-likelihood of this model by sampling the prior predictive log-likelihood of $\mathcal{D}$. For this, we use the same prior distribution $p_{M,T}(\theta)$ presented in 3.1 and sample $500$ predictions for $25$ data points, repeated with $5$ different splits of the dataset, to get samples of $p(\mathcal{D}|\alpha_{0})$
For in-context learning, we do not have access to $\varphi$ of the internal predictive model. However, since the language model provides us with predictions on targets given features directly, we can still approximate $p(\mathcal{D}|\alpha_{1})$ by repeated sampling of predictions. Here, we prompt the language model with our $100$ task descriptions and ask it to predict on $25$ data points from our dataset. We repeat this with $5$ different splits of the dataset to calculate the mean and standard deviation of the Bayes factor.
Figure 7 shows the distribution of the marginal likelihood for in-context learning, prior elicitation, and an uninformative prior along with the calculated Bayes factor (where reliably defined).
In the case of predicting Breast Cancer and Hypothyroidism, prior elicitation achieves a greater mean likelihood than both an uninformative prior and in-context learning. For these tasks, the Bayes factor and distribution of likelihoods suggest the use of prior elicitation, since it better describes the data (in terms of likelihood) and is cheaper to use for inference.
The Heart Disease prediction task favoured in-context learning, which provided a higher mean likelihood. For this task, the LLM returned an uninformative prior for the elicited distribution (Section 3.1), and used a prior with a smaller parameter range in-context (Section 3.3). In this case, the final choice of approach should depend on the setting, as using in-context learning for predictions might be too costly or unrealistic. Unlike the Bayes factor, the Bayesian Information Criterion (BIC) technique accounts for the computational complexity of classifiers (Section 2.4), but due to the large parameter sizes of an LLM and the availability of APIs, it could unfairly disfavour in-context learning. This approach may be more informative for some settings, and for Heart Disease prediction, would recommend prior elicitation.
For regression tasks, prior elicitation is preferred for $y=2x_{1}-x_{2}+x_{3}$, whilst for Diabetes, in-context learning produces greater log-likelihoods. However, here, both in-context learning and prior elicitation provide a high prior mean squared error and so a low marginal likelihood, making the Bayes factor unreliable to compute. Again, in the case of Diabetes prediction, the final model decision relies on the available resources.
Whether prior elicitation or in-context learning provides prior distributions with greater likelihoods depends on the task, with prior elicitation being the better choice over a uninformative prior for all datasets tested. However, when factoring in the methods’ abilities to estimate Bayesian inference and the cost of producing predictions, prior elicitation is likely the better option for the cases tested.
3.7 Choosing between priors
As seen in Appendices A.13.1 and A.13.2, the prior distributions elicited from language models do not always provide the best posterior performance. We consistently find that GPT-3.5-turbo and GPT-4-turbo provide better prior distributions than an uninformative prior; however, this may not be the case for all datasets or language models. Therefore, given prior distributions elicited from different language models, we would like to choose the best one for our predictive task. For this, given some collected data, we can use the Bayes factor (Section 2.4) as a model selection technique to decide between the prior distributions elicited from different language models. Alternatively, we could use other model selection techniques, such as the Bayesian Information Criterion. This is because we are performing the Bayesian inference ourselves, given the prior distributions. Furthermore, this method also allows us to compare the elicited distributions to the uninformative prior to ensure that the priors from our language models do not harm predictive performance. By using a temperature parameter, it is also possible to control the strength of the elicited prior, which can indicate our trust in the distribution.
4 Limitations and future research
In this work, we consider the language model as a domain expert to elicit prior distributions. These prior distributions need to be better than an uninformative prior to improve model performance and reduce data collection costs. A natural next step therefore is to question how these prior distributions compare to those elicited from domain experts, and whether language models provide an additional opportunity to reduce costs by supplementing expert elicited priors or by improving the expert elicitation framework. We began to explore how experts could work alongside an LLM by producing more detailed task descriptions (Section 3.2 and Appendix A.10), but other ways of integrating experts would be interesting to explore.
As priors elicited from language models are a product of their training, they could contain unwanted bias. An approach to avoid this would be to assign all features that could be considered sensitive to unwanted bias with an uninformative prior. However, prior distributions on other features might indirectly cause unwanted bias in a predictive model, and it is important to carefully think about whether these features should also be assigned an uninformative prior. Given an elicited prior distribution and observed examples, the Bayes factor can again be used for model selection to choose between combinations of an elicited prior with some or all features replaced with an uninformative prior. Additionally, by providing information in the task description, an expert could steer the language model away from unwanted bias in its provided distributions.
Finally, an interesting question is how we extend our results to neural networks. An immediate option is to ask a language model to suggest appropriate data augmentation techniques for a task, effectively reducing the cost of data collection by extending the current dataset. Another possible solution is to use the output of in-context predictions as additional features to be learnt from. Further exploration of methods in this direction would be beneficial.
5 Further results
In the Appendix, we supplement our findings with further results and discussions. Firstly, we present the complete prior elicitation and model selection process on a new dataset in Appendix A.5.5. Then we present additional results of prior elicitation and in-context learning (Appendix A.6, A.7, A.8), where we support the conclusions drawn in Sections 3. Next, since some of the datasets tested are in the public domain and might have appeared in the language model’s training, in Appendix A.9, we provide a discussion on whether the language model has memorised the datasets tested.
To further understand the language model’s behaviour to prompts, we study whether the language model is respecting our in-context model class specifications (Appendix A.3) and test how the detail of the task description affects the posterior performance of the elicited priors (Appendix A.10), as well as how experts can supplement the process in a controlled synthetic experiment. Further to this, we explore how the number of task descriptions affects the priors elicited (Appendix A.11).
Finally, to contextualise the value of prior elicitation from language models, we report the cost of performing prior elicitation (Appendix A.12) and evaluate the quality of priors elicited from other language models with different pricing tiers (Appendix A.13.1) as well as open-source models (Appendix A.13.2).
6 Conclusion
In this work, we presented a method for extracting task information from an LLM to build probabilistic linear models with fewer data points. By asking a language model to play the role of a domain expert, we were able to elicit prior distributions for predictive tasks, potentially reducing the cost of data collection and improving the accuracy of our models when we have access to few labelled examples (Section 3.1). As demonstrated in Appendix A.6.2, these prior distributions can significantly impact the delivery of care in healthcare settings, where label collection can be costly and time-consuming, but the need for accurate predictive models is great.
We find that the prior distributions elicited from GPT-3.5-turbo achieve better accuracy or mean squared error with fewer data points than an uninformative prior for a variety of tasks (Section 3.1). This performance improvement is most visible when few training examples were available, suggesting its use in environments where label collection is challenging. We also find that incorporating expert information into the task descriptions can enable further improvements in predictive accuracy, and that the language model updates its prior distributions as we would expect (Section 3.2 and Appendix A.10).
Since our approach to prior elicitation can also be seen as a direct replacement for in-context learning for datasets with numerical features, we present a model selection framework, based on the Bayes factor (Section 3.6). However, the Bayes factor assumes that the in-context learner is performing Bayesian inference. Therefore, we use maximum likelihood estimation and sampling distributions to approximate the internal model used during in-context learning and its prior and posterior distributions (Section 3.3). This allows us to inspect the truthfulness of the language model when we conduct prior elicitation (Section 3.4), as well as empirically question whether the language model is performing Bayesian inference on a given task (Section 3.5).
We find that prior elicitation is an improved alternative to in-context learning in our setting, as we can inspect priors, confidently perform Bayesian inference, and often achieve reduced error. Prior elicitation also extends the use of the language model’s prior knowledge of predictive tasks to previously infeasible situations, such as low-resource environments or where feature observations are frequent but labelled examples are few.
We have demonstrated that using language models for expert prior elicitation is a cost-effective method for improving the predictive performance of probabilistic linear models in environments with few labels and provides a competitive alternative to in-context learning.
Author Contributions
AC: Original idea conception, software, analysis of results, writing, editing, and reviewing, funding acquisition. RK: Original idea conception, supervision, analysis of results, reviewing. PB: Supervision, analysis of results, reviewing, funding acquisition.
Acknowledgments
This study is funded by the UK Dementia Research Institute (UKDRI) Care Research and Technology Centre funded by the Medical Research Council (MRC), Alzheimer’s Research UK, Alzheimer’s Society (grant number: UKDRI–7002), and the UKRI Engineering and Physical Sciences Research Council (EPSRC) PROTECT Project (grant number: EP/W031892/1). Infrastructure support for this research was provided by the NIHR Imperial Biomedical Research Centre (BRC) and the UKRI Medical Research Council (MRC). Further funding was provided by the Turing scheme. The funders were not involved in the study design, data collection, data analysis, or manuscript writing.
Appendix A Appendix
A.1 Availability of code
Within the supplementary materials, we provide code for generating all figures and results, as well as provide a pipeline for using our method on additional datasets.
Our code is implemented in python (v3.11.9) using pip (v24.0).
For our experiments, we use the Python packages:
  • blackjax (v1.2.3) [44]: For Monte Carlo sampling on the in-context prior samples.
  • jax (v0.4.31) [45]: For calculating some metrics and use with blackjax.
  • numpy (v1.26.4) [46]: For manipulation of arrays and computing mathematical equations.
  • openai (v1.51.2): For communicating with the OpenAI API.
  • pandas (v2.2.2) [47]: For manipulating the datasets and results.
  • pymc (v5.17.0) [48]: For Monte Carlo sampling on the elicited priors.
  • scikit-learn (v1.5.1) [49]: For loading and preprocessing the data.
  • transformers (v4.46.0): For loading and using Llama models.
A requirements file is provided within the supplementary code, which also includes some other utility packages required to run the scripts and load the data and results.
A.2 Formal details on extracting the internal prior and posterior
In this section, we provide a description of the method presented in Section 2.3.2 in more detail.
Given a language model $M$ and task $T$ that provides in-context predictions, if we can construct a function that allows us to sample a single parameterisation $\varphi\sim p_{M,T}(\varphi)$, then we can estimate $p_{M,T}(\varphi)$ using a sampling distribution. Following Murphy [32], we define an estimator:
$\varphi_{\text{MLE}}=\text{MLE}_{M,T}(\hat{\mathcal{D}})\\ =\arg\max_{\tilde{\varphi}}p_{M,T}(\hat{\mathcal{D}}|\tilde{\varphi})$ (5)
This corresponds to the maximum likelihood estimate (MLE) of $\varphi$ over some dataset $\hat{\mathcal{D}}$ given the task $T$. Now consider that $\hat{\mathcal{D}}$ is generated by a given $\varphi$ sampled from $\varphi\sim p_{M,T}(\varphi)$ and $X$. By sampling $x\sim\mathcal{U}[X]$ from a uniform distribution over $X$, we can create $S$ different datasets of the form:
$\hat{\mathcal{D}}^{(s)}=\{y_{n}\sim p_{M,T}(y_{n}|x_{n},\varphi):n=1:N,~{}x\sim\mathcal{U}[X]\}$ (6)
This provides us with the sampling distribution:
$\varphi_{\text{MLE}}^{(s)}\sim\text{MLE}_{M,T}(\hat{\mathcal{D}}^{(s)})\\ =\arg\max_{\tilde{\varphi}}p_{M,T}(\hat{\mathcal{D}}^{(s)}|\tilde{\varphi})$
With this, we can generate a sampling distribution using $\{\varphi_{\text{MLE}}^{(s)}\}_{s=0}^{s=S}$ that approximates $p_{M,T}(\varphi)$ as $S\rightarrow\infty$. Furthermore, by further incorporating the training dataset $\mathcal{D}$:
$\hat{\mathcal{D}}^{(s)}=\{y_{n}\sim p_{M,T}(y_{n}|x_{n},\varphi,\mathcal{D}):n=1:N,~{}x\sim\mathcal{U}[X]\}$
We get the sampling distribution:
$\varphi_{\text{MLE}}^{(s)}|\mathcal{D}\sim\text{MLE}_{M,T}(\hat{\mathcal{D}}^{(s)}|\mathcal{D})\\ =\arg\max_{\tilde{\varphi}}p_{M,T}(\hat{\mathcal{D}}^{(s)}|\tilde{\varphi},\mathcal{D})$
Allowing us to estimate $p_{M,T}(\varphi|\mathcal{D})$.
A.3 Does the language model use the model class we specify?
Figure 8: In-context predictions for the task: $y=2x_{1}-x_{2}+x_{3}$ compared to the true values. Here, feature values are generated using $X\sim\mathcal{U}(-5,5)$ and the targets are calculated using $y=2x_{1}-x_{2}+x_{3}$, which is specified fully in the in-context prompt. The language model is then asked to make predictions on $X$. This shows the true targets $y$ against those predicted by the language model, as well as the binned root mean squared error.
As part of comparing prior elicitation with in-context learning, we use maximum likelihood sampling to extract the internal parameter distributions (discussed in Section 3.3), which we use to test whether the language model is: (1) providing a similar prior distribution through elicitation to the one it is using in-context, and (2) empirically test whether the language model is approximating Bayesian inference. However, when extracting the internal parameters of the in-context model, we assume that the language model is applying the model class that we specify in the task description.
However, we find that for some datasets, the language model does not use the same model in context as the one we specify. For the synthetic data in particular, we specify the relationship as $y=2x_{1}-x_{2}+x_{3}$. Here, we can measure whether the language model is using the specified model in-context to predict the target by calculating the error between the true $y$ values and those returned by the language model’s in-context predictions. This is done by measuring the mean squared error between the $y_{n}$ given in Equation 6 and the true $\hat{y}_{n}=2x_{n,1}-x_{n,2}+x_{n,3}$ calculated using the $x_{n}$ in Equation 6.
In Figure 8 we show the true $\hat{y}_{n}$ against the in-context predictions $y_{n}$. Interestingly, even given the simple linear equation we ask the language model to mimic, we find that it struggles to provide good predictions in-context. This suggests that the language model is not always using the model class for in-context predictions that we specify in the task description, indicating that in its current state, in-context learning might be inadequate for predictive tasks with primarily numerical features. This figure also shows that the language model was hesitant to predict negative values and preferred positive predictions – a strange behaviour that clearly restricts its use as an emulator of a linear model.
However, for all other prior extraction tasks presented in Section 3.3, we find that the language model uses a linear model in-context, demonstrated by the low MSE loss values between the predictions of our MLE samples and the true in-context predictions, shown in Figure 12. This is often not the case when the language model is presented with demonstration samples, as in the posterior extraction presented in Figure 5, in which the language model uses a linear combination of features to predict the targets for the Heart Disease and Diabetes tasks only.
A.4 Energy distance and statistic
The energy statistic [33], which is based on the energy distance metric [50], is a method to measure the similarity of two probability distributions. Importantly, it allows us to test whether two $N$-dimensional probability distributions are equal.
The energy distance $D(X,Y)$, with $X$ and $Y$ as independent random vectors $\in\mathbb{R}^{n}$ is defined as:
$D(X,Y)^{2}=2\mathbb{E}[M_{XY}]-\mathbb{E}[M_{XX}]-\mathbb{E}[M_{YY}]$
Here, ${M_{XY}}_{i,j}=||x_{i}-y_{j}||_{2}$ is the pairwise distance matrix.
From this, we construct the energy statistic as follows:
$H(X,Y)=\frac{D(X,Y)^{2}}{2\mathbb{E}[M_{XY}]}$ (7)
This statistic takes values between 00 and $1$ and tests whether two statistical distributions are equal.
We provide an implementation of this distance metric and test statistic within our supplementary code (Appendix A.1).
A.5 Further experimental details
In all experiments, we use OpenAI’s GPT-3.5-turbo (specifically GPT-3.5-turbo-0125) with a temperature of $0.1$ and all other settings in the OpenAI API default  [^2] https://platform.openai.com/docs/api-reference/chat.
A.5.1 Dataset details
In this work, we evaluate our method on $6$ datasets; $1$ of which is synthetic, $1$ is privately collected, and $4$ are publicly accessible. This allows us to provide both reproducible experiments and test the language model’s ability to suggest a good prior on datasets not seen during training.
  • $y=2x_{1}-x_{2}+x_{3}$: To evaluate language models as tools for prior elicitation, we construct a synthetic dataset where we have control over the parameters of the dataset. Because of this, we are able to provide the language model with a complete description of the task. This dataset contains 3 features sampled from a normal distribution (mean $=0$ and standard deviation $=1$) and targets calculated using the equation $y=2x_{1}-x_{2}+x_{3}+\epsilon$, where $\epsilon\sim\mathcal{N}(0,0.05)$. When training and testing were done on these samples, no preprocessing was performed.
  • A private dataset collected as part of a study on Dementia care  [^3] https://www.imperial.ac.uk/uk-dri-care-research-technology/: This dataset contains $10$ features engineered from measurements of in-home activity and physiology. In this dataset, the labels corresponded to days with a clinical labelling of positive or negative urinary tract infection (UTI) which were calculated through the analysis of a provided urine sample in consultation with a healthcare professional. For this analysis, we used a subset of the data containing $78$ people and $371$ labelled days of UTI, with $270$ negatives and $101$ positives. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
  • The Heart Disease dataset [36, 37]: This dataset contains $10$ physiological features and targets with a positive or negative diagnosis of heart disease for $299$ people. Within this dataset, we have $161$ examples of people without heart disease and $138$ examples of people with heart disease, totalling $299$ data points. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data, except where features are categorical. In the categorical case, we perform no scaling.
  • The Hypothyroid dataset [39, 40]: This dataset, processed from the original, consists of $150$ examples of patients’ thyroid-related blood levels and whether they have a hypothyroidism diagnosis or not. This dataset contains $4$ features, with $75$ examples of a patient with hypothyroidism and $75$ examples of a patient without hypothyroidism. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
  • The Diabetes dataset [38]: After preprocessing, this dataset contains $442$ people with Diabetes and some physiological measurements. Along with some demographic information, the task is to predict a quantitative measure of their diabetes disease progression one year after this baseline. This dataset consisted of $10$ features and the targets were regression values with a maximum value of $10.0$, minimum value of $0.0$, and a mean value of $3.96$.
  • The Breast Cancer dataset [41]: This dataset is made up of $569$ data points with $30$ features computed from an image of a breast mass that are labelled as either malignant or benign. In total, this dataset contains $357$ examples of benign tumours and $212$ examples of malignant tumours. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
Where datasets are publicly available, we perform minimal processing for ease of reproducibility. In some cases, we replace the feature names with ones that are more descriptive (taken from the dataset descriptions also available publicly) to aid the language model’s comprehension of the data. All of these changes to the raw data are available within the supplementary code, which will automatically download and process the data when loading.
A.5.2 Evaluating prior elicitation
To evaluate the usefulness of the priors elicited from the language model (presented in Section 2.2), we started by comparing their posterior performance to an uninformative prior. To do this, we chose the predictive tasks described in Section 2.5: $y=2x_{1}-x_{2}+x_{3}$, Urinary Tract Infection (UTI), Heart Disease, Hypothyroidism, Diabetes, and Breast Cancer and calculated the posterior accuracy or mean squared error on a test set after observing varied numbers of examples. For each of the datasets, we selected $50\%$ of data points to form the test set and used the remaining data points to form a training set. This is done $10$ times to produce $10$ folds of random training and testing splits.
For the UTI task, we split the dataset so that each participant’s data is only in either the testing or training set, as some participants produced multiple examples. For the other classification tasks, the splits are stratified using the targets. To evaluate the posterior performance after observing a given number of data points, we randomly select that number of examples from the training set to train on and then test on the same test set as before.
After performing prior elicitation (Section 3.1), for each of the tasks, we have $100$ Gaussian priors for each feature in the dataset. We then build a mixture of the priors using Equation 1, where we weight the Gaussian distribution in each dimension and each task description separately.
We then construct a logistic or linear regression model, where we set the prior distribution over the bias term as a mixture of $100$ distributions of $\theta_{0}\sim\mathcal{N}(0,1)$. In the case of linear regression, we also use a noise term $\epsilon\sim\text{Half-Cauchy}(\beta=1)$, with a Half-Cauchy prior. In the uninformative case, we set $\theta\sim\mathcal{N}(0,1)$ for each feature.
We then use the No-U-Turn sampler [35] to sample from the posterior distribution over the linear model parameters, with $5$ chains and $5000$ samples per chain. In total, for each dataset, for each training and testing split, we have $25000$ posterior samples over the posterior.
To produce the results in Figures 2(a) and 9, we then sample the posterior predictive distribution over our test set and calculate the accuracy or mean squared error for each posterior parameter sample. To produce Figures 2(b) and 14, we also calculate the mean accuracy and the mean squared error for each of the test splits before plotting their mean and standard deviation.
As an example, for the synthetic task one of the task descriptions is as follows:
  • System role: You’re a linear regression predictor, estimating the target based on some input features. The known relationship is: ’target’ = 2 * ’feature 0’ - 1 * ’feature 1’ + 1 * ’feature 2’. Use this to predict the target value.
  • User role: I am a data scientist working with a dataset for a prediction task using feature values to predict a target. I would like to apply your model to predict the target for my samples. My dataset consists of these features: [’feature 0’, ’feature 1’, ’feature 2’]. All the values have been standardized using z-scores. The known relationship is: ’target’ = 2 * ’feature 0’ - 1 * ’feature 1’ + 1 * ’feature 2’. Based on the correlation between each feature and target, whether positive or negative, please guess the mean and standard deviation for a normal distribution prior for each feature. I need this for creating a linear regression model to predict the target. Provide your response as a JSON object with the feature names as keys, each containing a nested dictionary for the mean and standard deviation. A positive mean suggests positive correlation with the outcome, negative for negative correlation, and a smaller standard deviation indicates higher confidence. Only respond with JSON.
A.5.3 Extracting the in-context model’s hidden prior and posterior
To further study in-context learning, we used maximum likelihood sampling to extract approximate samples of the internal model’s prior and posterior using the method described in Section 2.3.2. For each of the tasks studied, we randomly generate $25$ feature values from $X\sim\mathcal{U}[-5,5]$ $5$ times and ask the language model to predict their regression label or positive classification probability based on the task description given and with the specified model class (in our case a linear or logistic regressor). This is done for each of the $100$ task descriptions that we have for the dataset. We then fit a linear regressor to the predicted regression labels or the classification logits  [^4] Calculated using the inverse logistic function: $\text{logits}=\ln p/(1-p)$ of the predicted probabilities, using maximum likelihood estimation. The parameters of this linear model are then taken as a single sample of the approximated in-context predictive model.
For each of these linear model samples, we calculate the regression labels or the classification logits using the same data given to the language model for in-context predictions. This allows us to calculate an error in our estimation of the in-context model, which depends on how well the in-context learner is approximating a linear model.
In total, for each task description, we collect $5$ samples of the in-context model parameters. Once this is repeated for all of the $100$ task descriptions, we have $500$ samples of the in-context model.
We can then measure, as in Section 3.4, whether the parameter samples are similar to those sampled from the elicited distributions. This is done by sampling $10000$ parameter values from the elicited distributions and calculating the energy statistic (Appendix A.4) between this sample and the extracted in-context sample and through visual inspection. By doing this, we test whether the language model is using the same distribution we elicit from it as the one it uses in-context. The results of this are presented in Table 1.
When extracting the posterior distribution over the in-context learner, we repeat the above process $5$ times, where in each repeat we provide the language model with $25$ demonstration points from the corresponding dataset. This allows us to empirically question whether the language model is approximating Bayesian inference when it updates the in-context prior distribution with the observed data points.
To do this, we start by fitting a Gaussian kernel density estimator (with a bandwidth factor of $0.25$) to our in-context model prior parameter samples. We then use the No-U-Turn sampler [35] with $1000$ adaption steps and the same demonstration examples given during in-context learning to sample posterior parameters. Here, we use $100$ chains of $10000$ samples to build an approximate posterior distribution.
We can then use the energy statistic to measure the difference between the extracted in-context posterior distribution and the Monte Carlo approximation of the posterior based on the extracted in-context prior distribution. The values presented in Table 2 show the result of this experiment.
A.5.4 Calculating Bayes factor
As part of our solution to decide whether to use prior elicitation or in-context learning for a given dataset, we calculate the Bayes factor (presented in Section 2.4).
For a single dataset, we start by randomly selecting $25$ data points, $5$ times, and consider the prior predictive distribution. For prior elicitation this involves sampling the prior predictive distribution (drawing $500$ samples for each split), and for in-context learning this involves prompting a language model for a prediction for each of the $100$ task descriptions (producing $100$ samples for each of the splits).
For both sets of prior predictive samples, we calculate their log-likelihood using either cross-entropy loss or mean squared error, depending on whether the task is classification or regression. For each dataset, for each test split, and for each method, we calculate the mean of this log-likelihood over the samples. With these values, we measure the difference between the value for the elicited prior predictions and the in-context predictions to obtain the log Bayes factor.
We then report the mean and standard deviation of the Bayes factor over the splits in Figure 7, which helps us to decide, given a dataset, which method is the most appropriate; assuming that in-context learning approximates Bayesian inference.
A.5.5 Given a new dataset
In all of our experiments, we asked the language model to mimic a linear model when predicting in-context, and to provide priors for a linear model when performing prior elicitation.
To apply the prior elicitation techniques to a new dataset, we recommend the following steps:
  1. 1.
    First, create a description of the dataset along with a description of the language model’s role. As part of this, provide descriptive features and target names. It will be helpful to split this task description into what the language model is expected to do (system role) and the task that the user is giving it (user role). Please see our collection of task descriptions in the supplementary code or the example given in Appendix A.5.2.
  2. 2.
    Ask a language model to rephrase the user role and system role many times. In our experiments, we asked the language model to rephrase both descriptions $10$ times for each.
  3. 3.
    Take the product of the system roles and user roles and provide each to the language model. For each of the unique system and user role combinations, record the mean and standard deviation of the Gaussian prior elicited from the language model.
  4. 4.
    Use these individual Gaussian priors to build a mixture of Gaussians as a prior distribution for a linear model.
  5. 5.
    Using the data points you have available from the predictive task, sample from the posterior predictive distribution to predict on unseen data.
Functions for doing this are provided in the supplementary code.
Now, to measure whether prior elicitation or in-context learning is better for a given task (assuming the language model is approximating Bayesian inference):
  1. 1.
    Firstly, sample the prior predictive distribution, with the elicited priors found above, on the data points already collected.
  2. 2.
    Then, to perform in-context learning, use a similar task description to above but with a different system and user role that specifies the language model should make predictions instead of providing prior distributions. With this new set of descriptions, again ask a language model to rephrase them.
  3. 3.
    To sample the prior predictive distribution using the product of these descriptions, ask a language model to make predictions on the data directly.
  4. 4.
    Measure the Bayes factor between the likelihood of the prior predictive distribution using prior elicitation and the in-context model on the already collected data.
  5. 5.
    Using the Bayes factor and by factoring in resource cost, decide whether in-context learning or prior elicitation is more appropriate.
To empirically check whether the language model is approximating Bayesian inference to produce its in-context predictions or to test whether it is being faithful to the priors we elicit, we can follow the below procedure:
  1. 1.
    Using the same task descriptions as above, provide the language model with randomly generated feature values (we used $25$ samples from $X\sim\mathcal{U}[-5,5]$) for each task description and ask it to make predictions in-context.
  2. 2.
    By additionally providing the seen data as demonstrations to the language model, again ask the language model to make predictions on randomly generated feature values for each task description.
  3. 3.
    With randomly generated feature values and in-context predictions, use maximum likelihood estimation to estimate the underlying linear model for each task description. For the classification case, this is done more easily if the language model is asked to provide the probability of a positive prediction rather than the prediction itself.
  4. 4.
    The resulting linear models approximated from the in-context predictions can now be used to approximate the prior and posterior distribution of the in-context model.
Using the above, we can now (1) empirically test whether the language model is being faithful to the prior distributions we elicit and (2) empirically test whether the language model is approximating Bayesian inference in-context.
For the former, we can use a measure of the difference in samples (in our case, the energy distance, Appendix A.4) to calculate the difference between the in-context model’s prior distribution and the elicited prior distribution. This will inform us of whether the language model is being truthful when it provides us with a prior on the parameters of a linear model, or whether the language model uses that prior knowledge when making predictions in-context. If the prior distribution elicited differs from the one that the language model is using in-context, then calculating the Bayes factor could be important to ensure that we select the most appropriate predictive model for the given task. If it is not clear whether the language model is approximating Bayesian inference in-context, we should then study the latter to ensure the Bayes factor is a suitable model selection tool.
For this, we can follow the below procedure:
  1. 1.
    Use a kernel density estimator (KDE) to approximate a smoother version of the prior distribution over the in-context model. For our experiments, we used a Gaussian kernel density estimator with a bandwidth factor of $0.25$.
  2. 2.
    Using the KDE and Monte Carlo methods, given the collected data, we sample a posterior distribution from the in-context learner’s prior.
  3. 3.
    Use a method to calculate the difference between two samples of a distribution to understand the differences between the posterior samples calculated through Monte Carlo sampling of the in-context prior and the samples of the in-context posterior.
This last procedure might indicate that our in-context learner is not approximating Bayesian inference and is not suitable for a given task, since it could be important to us that our predictive model is updating its prior knowledge in a reliable way.
A.6 Further prior elicitation results
Figure 9: Test accuracy of the posterior distribution over the model weights of a logistic and linear regression model for varied numbers of training data points. We evaluate the accuracy and mean squared error of different prior distributions over the weights of a logistic and linear regression model. Each of the Gaussians in the mixture is extracted from a language model and is tested against an uninformative prior $\mathcal{N}(\theta|0,1)$. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. These values are the result of $5000$ sampled parameters and $10$ random splits of the training and testing data. The green arrows point in the direction of improvement in the metric.
In Figure 9 we present the results of the same experiment as discussed in Section 3.1, applied to three other datasets discussed in Section 2.5.
For the synthetic task $y=2x_{1}-x_{2}+x_{3}$, the priors elicited by LLM allow for a posterior predictive distribution with a markedly reduced mean squared error over an uninformative prior and often produce less variation in the result. Due to the simple nature of the task however, after only a few training examples, both priors achieve similar posterior results.
In the case of Hypothyroid prediction, the LLM-elicited prior produces similarly accurate posterior parameter samples than the uninformative prior for smaller numbers of observations. However, as the number of observations increases, the LLM-elicited prior begins to produce more accurate posterior distributions.
For the Heart Disease dataset, we find that the prior elicited from the language model performed similarly to the uninformative prior. This is because the language model provided mean = 0 and std = 1 for a large number of priors elicited, suggesting that the language model was unsure about the prior distribution on the parameters for a model predicting heart disease. The parameter distributions elicited can be seen in Figure 12, which contextualises these results.
A.6.1 Priors elicited from the synthetic data
Figure 10 shows a histogram of parameter values from the GPT-3.5-turbo elicited prior for the task descriptions of $y=2x_{1}-x_{2}+x_{3}$. Here, the language model correctly provides prior distributions that are centred around the correct parameter values and offer a small standard deviation.
Figure 10: Elicited priors for the task $y=2x_{1}-x_{2}+x_{3}$. We sample $10000$ parameter values from the mixture generated from the Gaussian priors elicited from the language model when given the task description. This figure shows the histograms of those values in each feature dimension.
A.6.2 UTI accuracy with the label collection context
Figure 11: Posterior accuracy of the elicited priors on the UTI dataset considering label collection time. The average mean posterior accuracy and 95% confidence interval on the UTI test set after observing the given number of data points against the time it took for the data points to be collected in the study.
Since we have access to the underlying study that collected the UTI dataset, we can understand in more detail how the accuracy of the predictive model varies in time, as well as with the number of data points collected.
In Figure 11 we plot the time it took to collect the given number of labels tested in the experiments in Section 3.1 against the accuracy of a model using an uninformative prior and one elicited from a language model. This result illustrates the usefulness of prior elicitation, since we are able to achieve comparable (and ultimately better) accuracy with fewer data points, and therefore less data collection time.
Since UTI label collection is costly and slow, our UTI predictive model could have achieved greater accuracy significantly earlier in the study. This is paramount in a clinical environment, where an improved UTI detection model could have resulted in fewer unplanned hospital admissions.
It is important to note that the accuracy in Figure 11 is not achieved as a result of training on the earliest collected data points. This is because in our analysis we perform cross-validation and use a subset of the data from a later time period (Section A.5.1). Changes in the in-home devices were made during the study, such as the addition of a device to collect sleep information, which we use within the dataset evaluated in the main results. However, if all data collection devices had been used from the beginning of the study, we could expect the accuracy in time to resemble the figure presented here.
A.7 Is the language model being truthful when we elicit priors?
Figure 12: Distribution of the prior parameters under prior elicitation and those extracted from the internal model for the different predictive tasks. Here, we compare the prior distributions of the different predictive tasks, as defined by the internal model used in-context, and that which we elicit from the language model. Here, we draw the histogram separately in each dimension, but calculate the energy distance across all features combined.
In Section 3.4, we study how truthful the language model is when we elicit a prior distribution, compared to the prior it uses in its internal predictive model when performing in-context learning.
Figure 12 visualises the different prior parameter distributions for all of the tested datasets in each feature dimension. Notable are the large differences between the elicited prior distributions and the ones used when performing in-context learning, which supports the findings in Section 3.4. In the in-context case, prior distributions often have modes close to 00, suggesting that the internal model is not sure about the relationship between features and targets. The elicited priors often represent a stronger belief in the correlation between features and targets, with most histograms achieving modes greater than or less than 00.
Interestingly, on the Breast Cancer dataset, prior elicitation resulted in two distinct modes in the prior distribution for Feature 2 (“mean texture”), which could represent uncertainty in the direction of correlation.
Furthermore, the prior distributions on the parameters for the Heart Disease and Hypothyroid datasets have more variation in their values in comparison to that extracted from the language model’s internal model. This tells us that the elicited prior parameters represent more uncertainty in their value and are therefore more flexible when calculating the posterior on observed data.
A.8 Empirically, is the language model approximating Bayesian inference?
In this section, we present the differences between the posterior distribution extracted through maximum likelihood sampling of the LLM’s in-context internal model (Section 3.3) and one that is sampled through MC methods on the extracted in-context prior distribution.
Table 2: Difference between extracted posterior and MC on extracted prior. The energy statistic between the language model’s internal posterior extracted through MLE sampling and the posterior samples from applying Monte Carlo techniques to the language model’s internal prior. Values are $\in[0,1]$ with values closer to zero, indicating that the two distributions are more similar. Here, we show the mean and standard deviation of the energy statistic for the $5$ different training sets tested as observed data.
Heart Disease Hypothyroid Diabetes Breast Cancer
0.04 ± 0.01 0.06 ± 0.01 0.02 ± 0.00 0.01 ± 0.00
Figure 13: The different posterior samples Here, we compare the distribution of the posterior extracted from the language model’s in-context predictions against the distribution that we sample using Monte Carlo methods on the extracted prior distribution. Both posterior distributions are based on the same $25$ data point sample from the corresponding dataset. We draw the histogram separately in each dimension, but calculate the energy distance across all features combined.
Table 2 shows the mean and standard deviation of the energy distance between these distributions, for all of the datasets tested, including those that have a high approximation error when we estimate the in-context posterior. In addition, Figure 13 shows the distribution of these posterior parameters for visual inspection.
For the Breast Cancer dataset, these two distributions match closely and show a strong agreement in their parameter values, suggesting that for this dataset, in-context learning could be performing Bayesian inference. However, this should be considered alongside the results in Section 3.3, where we found that our approximations of the posterior distribution for the Breast Cancer task had a large error.
On the other hand, for the Hypothyroid predictive task, the parameter distributions differed significantly, with the in-context prior having a less varied distribution. This is reflected in the energy distance, which is the largest of all of the datasets tested.
These comparisons must be made with the approximation error in mind, presented in Section 3.3, which suggests that the posterior in-context distributions for the Breast Cancer and Hypothyroid tasks are not reliable.
A.9 Dataset memorisation
To better contextualise the use of in-context learning in the tested datasets, we will use the method proposed in Bordt et al. [42] to quantify the extent to which the language model already knows information about the datasets.
We start with the header completion test, which checks whether the language model has memorised the first few rows of the dataset. Here, the column names along with a given number of the first rows are given to the language model as a string, with the final of these rows truncated. The language model is then asked to generate the next $500$ tokens.
We then use the Levenshtein distance [51, 52], a string matching metric, to calculate the agreement between the true header completion and the one provided by the language model. The Levenshtein distance counts the number of single-character changes required to transform one string into another, only allowing for insertions, deletions, and substitutions. The value can take a minimum value of 00, suggesting that the strings are equal, and a maximum value of the longest length of the two strings.
To make the Levenshtein distance comparable across datasets, we calculate its normalised value by dividing the Levenshtein distance by the length of either the true string or the language model’s completion, whichever is longer. Because of this, the normalised Levenshtein distance reported is not formally a metric as it does not follow the triangle inequality. However, it allows us to compare the header completions across datasets more easily:
$\hat{L}(s_{1},s_{2})=\frac{\text{Levenshtein}(s_{1},s_{2})}{\text{min}\{|s_{1}|,|s_{2}|\}}$
Here, $\text{Levenshtein}(s_{1},s_{2})$ calculates the Levenshtein distance between two strings, and $\text{min}\{|s_{1}|,|s_{2}|\}$ returns the minimum length of the strings. This new value ranges between 00 and $1$, with 00 indicating a perfect match between the strings and $1$ implying that they were different at all positions.
Table 3: Header completion test results. Normalised Levenshtein distance between the correct header completion and the language model’s completion for each of the datasets tested. A value of 00 indicates that the true completion and language model’s completion were equal and a value of $1$ means that they were different at all positions.
Heart Disease Diabetes Hypothyroid Breast Cancer
0.00 0.21 0.22 0.00
The values in Table 3 demonstrate that the language model has completely memorised the first few rows of the Breast Cancer and Heart Disease datasets, demonstrated by a normalised Levenshtein distance value of 0. This could be because the first few rows of these datasets commonly appear within the language model’s training data. For the Diabetes and Hypothyroid datasets, we see some level of memorisation, with the language model being able to reproduce a large amount of the true header. As we might expect, when we perform the header test on the synthetic dataset generated using the relationship $y=2x_{1}-x_{2}+x_{3}$, we get a normalised Levenshtein distance of $0.7$, which shows a large difference between the true header string and the completion from the language model. The UTI task has not been tested, as it is a private dataset.
Table 4: Row completion test results. Normalised Levenshtein distance between the correct row completion and the language model’s completion for each of the datasets tested. A value of 00 indicates that the true completion and language model’s completion were equal and a value of $1$ means that they were different at all positions.
Heart Disease Diabetes Hypothyroid Breast Cancer
0.33 ± 0.05 0.38 ± 0.05 0.20 ± 0.05 0.43 ± 0.13
We supplement these results with those of Table 4, which shows the results of the row completion test. Here we calculate the normalised Levenshtein distance of the true completions and the language model’s completion of a random row from the dataset. In these tests, the language model is provided with $10$ rows (in order from the original dataset) and is asked to complete the following row. The values in Table 4 represent the mean and standard deviation of the test using $25$ random rows. Here, in contrast to the header test, we see that the language model has not memorised all rows from the dataset. This is because the language model’s completion differs from the true row completion in all cases.
We likely see these results because when the datasets appear within the language model’s training data, their column names and initial rows are usually given, but not the entire dataset.
A.9.1 How does this affect our results?
Within our experiments, we normalise the training and testing data, as well as sample random observations from the dataset. Because of this, observations predicted during in-context learning will likely not have been memorised, since they are normalised based on a training data subset and are selected from random rows of the dataset. As shown, although the language model can complete the first few rows of the datasets with some success, it is unable to complete random rows (Table 4). Therefore, we hypothesise that any memorisation of the tested datasets has little impact on the in-context predictions.
Further, since for prior elicitation we only prompt the language model for a prior distribution of a predictive model given the feature names and a description of the dataset, the memorisation of rows from the data is unlikely to have had an impact on these results. In addition, we see that elicited priors provide improved predictions over the uninformative prior for the private and synthetic datasets, as well as the public datasets.
A.10 Task descriptions with more information
Figure 14: Mean squared error using the elicited priors for the task: $y=2x_{1}-x_{2}+x_{3}$ where we have used increasingly more detailed task descriptions. Here, feature values are generated using $X\sim\mathcal{U}(-5,5)$ and the targets are calculated using $y=2x_{1}-x_{2}+x_{3}$. Within each task description we provide the language model with increasingly more information about the task, as to test how the task description effects the priors elicited. Complete descriptions of the task descriptions given here are provided in Section A.10. The lines depict the average mean posterior mean squared error (MSE) over the 5 dataset splits, whilst the error bars represent the 95% confidence interval.
To understand how the detail of the task description affects the performance of the priors elicited from the language model, we studied the synthetic task, $y=2x_{1}-x_{2}+x_{3}$, and described the relationship between features and targets in varied detail. We additionally test an uninformative prior to contextualise the results. For each of the levels of detail shown in Figure 14, we included the following text in the task description that was provided to the language model:
  • Linear relationship:
    “The target is linear in features”
  • One feature relationship:
    “The target is a linear combination of the features and that when ‘feature 0’ increases by 1, the target increases by 2”
  • Two feature relationships:
    “The target is a linear combination of the features and that when ‘feature 0’ increases by 1, the target increases by 2, and when ‘feature 1’ increases by 1, the target decreases by 1”
  • Three feature relationships:
    “The target is a linear combination of the features and that when ‘feature 0’ increases by 1, the target increases by 2, and when ‘feature 1’ increases by 1, the target decreases by 1, and when ‘feature 2’ increases by 1, the target increases by 1”
  • Full equation:
    “The ‘target’ = 2 * ‘feature 0’ - 1 * ‘feature 1’ + 1 * ‘feature 2’”
For each prior elicited using these levels of detail, in Figure 14, the corresponding $\log_{10}\text{MSE}$ on the test set can be seen for a linear regressor after observing the given number of data points. This figure illustrates the effects of varied descriptions and shows that the improvements made to the MSE diminish as the task description contains more detail. The biggest improvement came from providing the language model with a single feature relationship – enabling an MSE that is two orders of magnitude better. This suggests that it is most helpful to provide more detail when the language model initially knows very little about a task.
Providing the language model with improved levels of detail allows it to produce prior distributions that enable our linear regression model to achieve a given error rate with fewer data points. As we might expect, the more detail provided in the task description, the better our prior distributions are.
However, what is unexpected is that for this task, using an uninformative prior ($\theta\sim\mathcal{N}(\theta|0,1)$) provides better posterior performance than specifying to the language model that the relationship between the target and features is linear (Figure 14). This is explained by inspecting the prior distribution elicited when given this minimal information, which shows that the language model responded with a similar mean and standard deviation for a Gaussian for each of the $100$ variations of the prompts:
  • Feature 0: mean: 0.5 ± 0.11, std: 0.12 ± 0.05
  • Feature 1: mean: -0.32 ± 0.08, std: 0.08 ± 0.06
  • Feature 2: mean: 0.56 ± 0.26, std: 0.14 ± 0.09
This demonstrates the sometimes unexpected behaviour of language models, which here produced Gaussians for the prior distribution with small standard deviations (reflecting a high confidence in the distribution), and means that imply the language model knew the underlying equation defining the relationship between the target and features. We might have expected that after only being told the relationship was linear, the language model would produce Gaussians with 00 mean and standard deviation $1$, reflecting the uncertainty in the exact relationship between each feature and the target.
Interestingly, although “Three feature relationships” and “Full equation” contain the same information written differently, the language model provided better prior distributions when given the full equation. Fully specifying the equation is not realistic in a real-world setting where we are not sure about the exact relationship between features and targets, but this provides an example of how we can make the task description as information dense as possible.
Figure 15: Parameter distribution as further feature information is included. These histograms show the distribution of parameter values for the features as more expert information is provided. In each row, we provide information in the task description of an additional feature, with the green arrow indicating the effect this has on the parameter distribution. In particular, in the first row we provide information about the relationship between the target and the first feature, in the second row we provide information about the second feature, and the in the third row we provide information about the third feature.
An expert interacting with the language model might be able to provide further information about a task that improves the elicited priors. In Figure 15 we show the parameter distributions elicited from the language model for the different descriptions we provide. This figure shows that, as we provide additional information about each feature, the language model responds by providing a prior distribution that reflects its updated knowledge of the task. This demonstrates the potential of using an expert to construct more detailed task descriptions, which the language model can translate into parameter distributions. This experiment provides a controlled setting that supplements the results in Section 3.2.
A.11 How many ways should we describe a task?
Figure 16: Measuring how similar the prior distributions are with fewer task descriptions. We reduce the number of task descriptions and calculate the energy statistic between $10000$ samples from the prior elicited from the language model using the reduced set of descriptions and the full $100$ task descriptions. We repeat each experiment $10$ times and show the mean and standard deviation.
In this section, we look at how many of the task descriptions are required to be able to get a similar prior distribution to that elicited within our experiments, using $100$ task descriptions.
Figure 16 shows the energy statistic (Appendix A.4) of samples from the prior distribution elicited using fewer task descriptions compared to samples from the prior distribution elicited using $100$ task descriptions (from the same language model). In particular, we compare a subset of task descriptions with the full $100$ task descriptions.
We see that for most tasks, a large number of task descriptions is required, since reducing the number of descriptions from $100$ to $49$ produced prior distributions that are significantly different for all datasets tested. Using $81$ task descriptions produces prior distributions that are more similar but still show differences for some datasets (for example, $y=2x_{1}-x_{2}+x_{3}$ and Breast Cancer).
It is interesting to see that Breast Cancer benefitted most from increasing the number of task descriptions, since this dataset contains the largest number of features. It suggests that more task descriptions are required for those datasets with high-dimensional data. This makes intuitive sense since we would expect to require a larger mixture of Gaussian components to better describe the prior over for a higher-dimensional dataset.
This figure shows that as the number of task descriptions increases, we see smaller changes in the energy distance and reduced variability in the samples. This suggests that increasing the number of task descriptions provides diminishing returns.
Further work could allow the number of task descriptions to be a function of the number of features in the dataset.
A.12 Cost of elicitation
The priors elicited from GPT-3.5-turbo incur a cost, as we use OpenAI’s API. To elicit all of the priors used in the experiments in Section 3.1, we were charged \$$\$0.41$ (on average \$$\$0.068$ per dataset). However, we did not make use of the batch API and instead made each API call sequentially. All experiments in this work could be implemented using the batch API since no new API call depends on a previous API result, which may significantly reduce costs  [^5] https://openai.com/api/pricing/.
A.13 Other language models
In this section, we will compare the results presented in the main paper (on GPT-3.5-turbo) to some other commonly used language models. We start by exploring the performance of some other OpenAI models  [^6] https://platform.openai.com/docs/models, before considering some of the open source Llama models  [^7] https://huggingface.co/meta-llama [15, 16, 17].
A.13.1 Other OpenAI models
Figure 17: Test accuracy of the posterior distribution for different OpenAI language models. Here we show the posterior results when using priors elicited from two other OpenAI language models and an uninformative prior $\mathcal{N}(\theta|0,1)$. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. The values are the result of $5000$ sampled parameters for each of $5$ chains, and $10$ random splits of the training and testing data. For the regression tasks, we report the mean squared error on log-scale. This figure shows the line plot of the average mean posterior accuracy or mean squared error over the $10$ splits, with the error bars representing the 95% confidence interval. The green arrows point in the direction of improvement in the metric.
In this section we compare the posterior results of priors elicited from GPT-3.5-turbo to GPT-4-turbo which incurs a significantly higher cost, and GPT-4o-mini which provides a less expensive option.
Specifically, to elicit all priors presented here, GPT-4o-mini cost \$$\$0.32$ (on average \$$\$0.053$ per dataset), whilst GPT-4-turbo cost \$$\$7.56$ (on average \$$\$1.26$ per dataset). This is compared to the cost of \$$\$0.41$ (on average \$$\$0.068$ per dataset) for GPT-3.5-turbo (presented in A.12). As a rule, the more features a dataset contains or the longer the task description, the more expensive priors are to elicit since the number of input and output tokens is greater.
Figure 17 shows that in most cases, GPT-4-turbo provides priors that are significantly better than GPT-3.5-turbo but falls behind in the synthetic and Hypothyroid prediction tasks. On the Breast Cancer, UTI, and Heart Disease tasks, the priors elicited from GPT-4-turbo allow for a greatly improved posterior accuracy, and unlike GPT-3.5-turbo, provide an informative prior for Heart Disease. In addition, on the Diabetes predictive task, GPT-4-turbo achieves similar mean squared error to GPT-3.5-turbo, with a small improvement in performance when only a few training data points are available.
On the other hand, GPT-4o-mini often produces prior distributions that are equal to or worse than GPT-3.5-turbo in terms of their posterior performance in all tasks except one. For the synthetic, Diabetes, and Heart Disease prediction tasks, GPT-4o-mini provides prior distributions that achieve similar or worse performance to the uninformative prior. However, for the UTI and Breast Cancer predictive tasks they achieve similar posterior accuracy to GPT-3.5-turbo at a lower price point. Also, on the Hypothyroid dataset GPT-4o-mini provides prior distributions that achieve greater accuracy than those from GPT-3.5-turbo and GPT-4-turbo for all training sizes, whilst requiring significantly less cost.
A.13.2 Open source language models
Figure 18: Test accuracy of the posterior distribution for Llama language models. Here, we show the results when using priors elicited from two of the Llama language models and an uninformative prior $\mathcal{N}(\theta|0,1)$ on a linear model. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. The values are the result of $5000$ sampled parameters for each of $5$ chains, and $10$ random splits of the training and testing data. For the regression tasks, we report the mean squared error on log-scale. This figure shows the line plot of the average mean posterior accuracy or mean squared error over the $10$ splits, with the error bars representing the 95% confidence interval. The green arrows point in the direction of improvement in the metric.
Here, we explore the posterior performance of prior distributions elicited from three Llama [15, 16, 17] models and one Qwen model [53, 54]:
In contrast to the OpenAI models in Appendix A.13.1, these models are publicly available and open source. This means that they can be used without the need for connecting to an online API, and when locally hosted, all data and inference can be performed on device, enabling the use of our framework within a closed and secure environment.
The experiments presented in this section use the same task descriptions as given to GPT-3.5-turbo, except that because these smaller models sometimes fail to return a JSON structure, we suffix the prompt with an additional instruction to return a JSON. When sampling the posterior with the elicited priors, we use the same methods and options as given in A.5.2.
Figure 18 illustrates the inconsistent posterior accuracy of the priors elicited from the Llama family of tested language models. These models are significantly smaller in capacity to GPT-3.5-turbo, justifying their inconsistent elicited priors, but in some of the datasets, they still produce the greatest posterior performance.
Specifically for Breast Cancer, the 3B parameter Llama model provides an intermediate accuracy between an uninformative prior and those elicited from GPT-3.5-turbo, and for the synthetic and Hypothyroid tasks, the 8B parameter Llama provides priors of equal quality to GPT-3.5-turbo. However, for the UTI and Heart Disease datasets, both the 3B and 8B Llama models provide prior distributions that are of worse quality than even the uninformative prior. It is possibly not surprising that the Llama-3.2-3B-Instruct provided priors that were equal or worse than Llama-3.1-8B-Instruct in all cases except for predicting Breast Cancer since these models are significantly smaller in size.
Llama-3.1-70B-Instruct-BF16 is more competitive, sometimes providing priors that allow for better posterior distributions than GPT-3.5-turbo, but on other occasions, providing priors with similar posterior performance to the uninformative prior. On the UTI and Heart Disease datasets, the prior distributions elicited from Llama-3.1-70B-Instruct-BF16 provide better posterior accuracy than those from GPT-3.5-turbo. Surprisingly, on Heart Disease prediction in particular, priors from the 70B Llama model provided a posterior accuracy greater than GPT-4-turbo. Additionally, when predicting Breast Cancer, priors form Llama-3.1-70B-Instruct-BF16 achieved the same posterior accuracy as GPT-3.5-turbo for all of the number of training examples tested. These three datasets suggest that Llama-3.1-70B-Instruct-BF16 can be a useful model for eliciting prior distributions for a predictive task. However, when eliciting priors for the Diabetes and synthetic tasks, we found that Llama-3.1-70B-Instruct-BF16 provided distributions that were not competitive with GPT-3.5-turbo.
Qwen-2.5-14B-Instruct is similarly inconsistent in the posterior performance that the priors provide. For Heart Disease and Hypothyroid prediction, it achieves a greater posterior accuracy than all models in Figure 18, except for the 70B Llama model, Llama-3.1-70B-Instruct-BF16. For the UTI task, after seeing $10$ demonstrations, it performs comparably to Llama-3.1-70B-Instruct-BF16, and better than GPT-3.5-turbo. However, on the Breast Cancer and synthetic tasks, it performs worse than GPT-3.5-turbo, but is likely smaller in capacity. Overall, Qwen-2.5-14B-Instruct provides inconsistent performance across tasks, but has the benefit of performing similarly to Llama-3.1-70B-Instruct-BF16 on many tasks with a significantly smaller compute requirement.
It is important, therefore, given an elicited prior distribution, to calculate the prior likelihood with some already collected data to understand whether that distribution is helpful for constructing a predictive model.